### v1

# Dynamics of the sine-Gordon equation on tadpole graphs

Abstract
This work studies the dynamics of solutions to the sine-Gordon equation posed on a tadpole graph $\mathcal{G}$ and endowed with boundary conditions at the vertex of $\delta$-type. The latter generalize conditions of Neumann-Kirchhoff type. The purpose of this analysis is to establish an instability result for a certain family of stationary solutions known as single-lobe kink state profiles, which consist of a periodic, symmetric, concave stationary solution in the finite (periodic) lasso of the tadpole, coupled with a decaying kink at the infinite edge of the graph. It is proved that such stationary profile solutions are linearly (and nonlinearly) unstable under the flow of the sine-Gordon model on the graph. The extension theory of symmetric operators, Sturm-Liouville oscillation results and analytic perturbation theory of operators are fundamental ingredients in the stability analysis. The local well-posedness of the sine-Gordon model in an appropriate energy space is also established. The theory developed in this investigation constitutes the first stability result of stationary solutions to the sine-Gordon equation on a tadpole graph.
1 Department of Mathematics, IME-USP
Rua do Matão 1010, Cidade Universitária, CEP 05508-090, São Paulo, SP (Brazil)
angulo@ime.usp.br
2 Instituto de Investigaciones en Matemáticas Aplicadas y en Sistemas,
Universidad Nacional Autónoma de México, Circuito Escolar s/n,
Ciudad Universitaria, C.P. 04510 Cd. de México (Mexico)
plaza@aries.iimas.unam.mx
Mathematics Subject Classification (2020). Primary 35Q51, 35J61, 35B35; Secondary 47E05.
Key words. sine-Gordon model, metric graphs, kink and anti-kink solutions, $\delta$-type interaction, analytic perturbation theory, extension theory, instability.
1 Introduction
Nonlinear evolution models on quantum graphs emerge as useful tools for studying wave propagation, particularly in quasi-one-dimensional systems such as meso- or nanoscale structures that resemble a thin neighborhood around a graph. These models offer a practical way to explore various physical phenomena in both theoretical and experimental contexts. The flexibility in configuring the graph’s geometry enables the generation of diverse dynamics. Examples of bifurcated systems include Josephson junction structures, networks of planar optical waveguides and fibers, branched structures like DNA, blood pressure waves in large arteries, nerve impulses in complex neuron arrays, conducting polymers, and Bose-Einstein condensation (see, e.g., [20, 23, 24, 26, 27, 34, 36, 41, 48] and references therein).
Recently, nonlinear models on metric graphs, such as the nonlinear Schrödinger equation, the sine-Gordon model, and the Korteweg-de Vries equation have garnered considerable attention (for an abridged list of related works the reader is referred to [1, 12, 11, 8, 9, 14, 15, 16, 2, 27, 41]). Mathematically, an evolution model on a graph is equivalent to a system of partial differential equations (PDEs) defined along the edges (viewed as intervals), with the coupling determined solely by the boundary conditions at the vertices—this is referred to as the “topology of the graph”. These boundary conditions govern the network’s dynamics through group interactions. In particular, the study of soliton transport has been particularly prominent in this field, as solitons and other nonlinear waves in branched systems provide valuable insights into the behavior of these models.
Addressing these challenges is difficult due to the complexity of both the equations of motion and the graph topology. A particular issue complicating the analysis is the behavior at the vertices, where a soliton profile traveling along one edge may experience complex phenomena like reflection and radiation emission. This makes it hard to track how energy propagates through the network. Consequently, studying soliton propagation in networks presents significant obstacles. The mechanisms governing the existence and stability (orbital or spectral) of soliton profiles remain uncertain for many types of graphs and models. This project intends to contribute with the exploration of these topics in the particular case of the sine-Gordon equation,
$\partial_{t}^{2}\textbf{U}-\textbf{c}^{2}\Delta\textbf{U}+\sin\textbf{U}=0.$ (1.1)
We recall that a metric graph $\mathcal{G}$ is a structure represented by a finite number of vertices $V=\{\nu_{i}\}$ and a set of adjacent edges at the vertices $E=\{e_{j}\}$ (for further details, see [20]). Each edge $e_{j}$ can be identified with a finite or infinite interval of the real line, $I_{e}$. Thus, the edges of $\mathcal{G}$ are not merely abstract relations between vertices but can be viewed as physical “wires” or “networks” connecting them. The notation $e\in E$ will be used to indicate that $e$ is an edge of $\mathcal{G}$. This identification introduces the coordinate $x_{e}$ along the edge $e$. A quantum graph is a metric graph $\mathcal{G}$ equipped with a differential (or pseudo-differential) operator acting on functions on the graph.
One of the main graphs of interest to us here is the so-called tadpole graph (also known as lasso graph [26]), a graph composed of a circle and a infinite half-line attached to a common vertex $\nu=L$. If the circle is identified with the interval $[-L,L]$ and the half-line with $[L,\infty)$, we obtain a particular metric graph structure that we will denote again by $\mathcal{G}$, represented by $E=\{[-L,L],[L,\infty)\}$ and $V=\{L\}$ (see Figure 1). We use the subscripts $1$ and $2$ to refer to the edges $e_{1}=[-L,L]$ and $e_{2}=[L,\infty)$.
Figure 1: Tadpole graph $\mathcal{G}$ with vertex at $\nu=L$.
A wave function U defined on the tadpole graph $\mathcal{G}$ will be understood as an pair of functions $\textbf{U}=(u_{1},u_{2})$, where $u_{1}$ is defined on $E_{1}=[-L,L]$ and $u_{2}$ on $E_{2}=[L,\infty)$. Thus, we define the action of the Laplacian operator $\Delta$ on the tadpole graph $\mathcal{G}$ by
$-\Delta:(u_{e})_{e\in E}\to(-u^{\prime\prime}_{e})_{e\in E}.$ (1.2)
Thus, from (1.1), for $\mathbf{U}(x_{e},t)=(u_{1}(x_{e},t),u_{2}(x_{e},t))$ and the nonlinearity $\sin\mathbf{U}$ acting componentwise, i.e., $\sin\mathbb{U}(x_{e},t)=(\sin u_{1}(x_{e},t),\sin u_{2}(x_{e},t))$, we obtain that for $\textbf{c}^{2}=(c_{1}^{2},c_{2}^{2})$ the model in (1.1) can be written in the following vectorial form on a tadpole graph $\mathcal{G}$,
$\partial_{t}^{2}u_{j}-c_{j}^{2}\partial_{x}^{2}u_{j}+\sin u_{j}=0,\qquad x\in E_{j},\quad c_{j}>0,\quad j=1,2.$ (1.3)
Posing the sine-Gordon equation on a metric graph comes out naturally from practical applications. For example, since the phase-difference in a long (infinite) Josephson junction obeys the classical sine-Gordon model on the line (cf. [18, 32, 50]), the coupling of two or more Josephson junctions forming a network can be effectively modeled by the vectorial sine-Gordon model in (1.1) on a graph. The sine-Gordon equation was first conceived on a $\mathcal{Y}$-shaped Josephson junction by Nakajima et al. [40, 39] as a prototype for logic circuits. The authors consider three long (semi-infinite) Josephson junctions coupled at one single common vertex, a structure known as a tricrystal junction. There exists two main types of $\mathcal{Y}$-junctions. A $\mathcal{Y}$-junction of the first type (or type I) consists of one incoming (or parent) edge, $E_{1}=(-\infty,0)$, meeting at one single vertex at the origin, $\nu=0$, with other two outgoing (children) edges, $E_{j}=(0,\infty)$, $j=2,3$. The second type (or $\mathcal{Y}$-junction of type II) resembles more a starred structure and consists of three identical edges of the form $E_{j}=(0,\infty)$, $1\leqq j\leqq 3$. Junctions of type I are more common in unidirectional fluid flow models (see, for example, [21]), whereas graphs of type I or II are indistinctively used to describe Josephson tricrystal junctions; see, for instance, [29, 49] (type I), or [50, 33, 46] (type II). Recently, Angulo and Plaza [14, 15, 16] studied the sine-Gordon model on $\mathcal{Y}$-junctions of type I and II by obtaining an comprehensive study of the local well-posedness for the Cauchy problem and the linear (and nonlinear) instability results of static, kink and kink/anti-kink soliton profile solutions. The established theory considers boundary conditions at the graph-vertex of either $\delta$-interaction type (expressing continuity of the wave functions plus a Kirchhoff-type rule expressing conservation of the current low, the Kirchhoff law for electric currents) or $\delta^{\prime}$-interaction type (expressing continuity of the derivatives of the wave functions plus a Kirchhoff-type rule for the self-induced magnetic flux).
The goal of this paper is to shed new light on the study of the dynamics of the sine-Gordon vectorial model in (1.1) in the case of a tadpole graph. Hence, we will consider the sine-Gordon model (1.3) as a first order system on the bounded interval and on the half-line, respectively,
[…]
with $u_{j}=u_{j}(x,t)$, $v_{j}=v_{j}(x,t)$ and $u_{1},v_{1}:[-L,L]\to\mathbb{R}$, $u_{2},v_{2}:[L,\infty)\to\mathbb{R}$. Here, we are interested in the dynamics generated by the flow of the sine-Gordon on the tadpole around solutions of static or stationary type,
$u_{j}(x,t)=\phi_{j}(x),\qquad v_{j}(x,t)=0,$ (1.5)
for all $j=1,2$, and $x\in E_{j}$, $t>0$, where each of the profile functions $\phi_{j}$ satisfies the equation
$-c_{j}^{2}\phi^{\prime\prime}_{j}+\sin\phi_{j}=0,$ (1.6)
on each edge $E_{j}$ and for all $j$, as well as the boundary conditions at the vertex $\nu=0$ being the so-called $\delta$-coupling or generalized Neumann-Kirchhoff boundary conditions, namely, $(\phi_{1},\phi_{2})\in D_{Z}$
$D_{Z}=\{(f,g)\in H^{2}(-L,L)\times H^{2}(L,\infty):f(L)=f(-L)=g(L),\,f^{\prime}(L)-f^{\prime}(-L)=g^{\prime}(L)+Zg(L)\},$ (1.7)
where for any $n\geqq 0$,
$H^{n}(\mathcal{G})=H^{n}(-L,L)\oplus H^{n}(L,\infty).$
The parameter $Z\in\mathbb{R}$ is a coupling constant between the disconnected loop and the half-line. The values of $f$, $g$ and their derivatives at the vertex (and at $-L$) are understood as the appropriate one-sided limits (observe that, as $(f,g)\in H^{2}(-L,L)\times H^{2}(L,\infty)$, $f^{\prime}(\pm L)$ and $g^{\prime}(L+)$ are well defined). We note that $(-\Delta,D_{Z})_{Z\in\mathbb{R}}$ represents a one-parameter family of self-adjoint operators on a tadpole graph (see Theorem A.4 in the Appendix).
The delicate point about the existence of stationary solutions solving (1.6) is given by the component $\phi_{1}$ on $[-L,L]$. The profile $\phi_{2}$ on $[L,\infty)$ is given by the well-known kink-soliton profile on the full real line modulo translation (see [25, 47]),
$\phi_{2}(x)\equiv\phi_{2,a}(x)=4\arctan\big{[}e^{-\frac{1}{c_{2}}(x-L+a)}\big{]},\quad\;\;x\in[L,\infty),$ (1.8)
where we are interested for the case $c_{2}>0$ ($\lim_{x\to\infty}\phi_{2}(x)=0$) and the shift parameter $a\in\mathbb{R}$ will be determined by the condition $(\phi_{1},\phi_{2})\in D_{Z}$ (see Figure 2).
Figure 2: Graph of the kink-soliton profile (in solid blue line) for $x\in[L,\infty)$ for different values of $a$. Here $L=2$ and $c_{2}=1$ (color online).
Among all stationary solutions for (1.6) (specifically on the lasso, where we can see the existence of infinite profiles with many lobes), we are particularly interested in the so-called single-lobe kink states, an example of which is shown on Figure 3. More precisely, we have the following definition.
Figure 3: An example of the graph a single-lobe kink state profile for the sine-Gordon model according to Definition 1.1.
Definition 1.1.
The stationary profile solution $\Theta=(\phi_{1},\phi_{2})\in D_{Z}$ is said to be a single-lobe kink state for (1.6) if every component is positive on each edge of the tadpole graph $\mathcal{G}$, the maximum of $\phi_{1}$ is achieved at a single internal point, symmetric on $[-L,L]$ and monotonically decreasing on $[0,L]$. Moreover, $\phi_{2}$ is strictly decreasing on $[L,\infty)$.
The existence and dynamics of positive single-lobe kink states for the sine-Gordon model on a tadpole graph have not been addressed in the literature and, consequently, one of the purposes of this article is to take the first steps towards a better understanding of the dynamics of these profiles by the flow of the sine-Gordon model. In particular, in Proposition 4.1 below, we show that there are single-lobe kink states $(\phi_{1},\phi_{2})\in D_{Z}$ provided that $Z\in(-\infty,\frac{2}{\pi c_{2}})$.
For the convenience of the reader we give a brief description of our main results. In Propositions 4.2 and 4.4 below, we establish the profiles of a two-families of single-lobe states for (1.6) on $[-L,L]$ depending of the signal of the kink-shift, $a<0$ or $a>0$ (but no necessarily being the first component for a single-lobe kink state $(\phi_{1},\phi_{2})\in D_{Z}$). In Proposition 4.2 ($a<0$), this family of solutions $\phi_{1}\equiv\phi_{1,k}$ for equation in (1.6) are a piece of subluminal periodic type solution for the sine-Gordon with a single-lobe profile on $[-L,L]$ given by,
$\phi_{1,k}(x)=2\pi-\arccos\Big{[}-1+2k^{2}\operatorname{sn}^{2}\Big{(}\frac{x}{c_{1}}+K(k);k\Big{)}\Big{]},\quad x\in[-L,L],$ (1.9)
where $\operatorname{sn}(\cdot;k)$ is the Jacobian elliptic function of snoidal-type and the elliptic modulus $k$ satisfying $K(k)>\frac{L}{c_{1}}$. More precisely, for $\frac{L}{c_{1}}\leqq\frac{\pi}{2}$, we have $k\in(0,1)$, and for $\frac{L}{c_{1}}>\frac{\pi}{2}$, we need to have $k\in(k_{0},1)$, with $K(k_{0})=\frac{L}{c_{1}}$, by properties of the strictly increasing Legendre’s complete elliptic integral of the first type $K$ (see, e.g., Byrd and Friedman [22]). The profile $\phi_{1,k}$ satisfies $\phi^{\prime\prime}_{1,k}(x)<0$ for all $x\in[-L,L]$ (see Figure 4). Moreover, In Proposition 4.7 below we give a complete restrictions on the parameters $L,c_{1},c_{2},k$ and $Z$ for which there are (or there are not) single-lobe kink states $(\phi_{1,k},\phi_{2,a(k)})\in D_{Z}$.
Figure 4: Single-lobe solution $\phi_{1,k}$ in (1.9) with $L=\pi$, $c_{1}=1$ and $k^{2}=0.98$.
The second goal of the present analysis is associated with the dynamics of the sine-Gordon flow around the kink profiles obtained in Proposition 4.7; more specifically, we are interested on its stability properties. The stability of these static configurations is an important property from both the mathematical and the physical points of view. Stability can predict whether a particular state can be observed in experiments or not. Unstable configurations are rapidly dominated by dispersion, drift, or by other interactions depending on the dynamics, and they are practically undetectable in applications.
In the forthcoming stability analysis, the family of linearized operators around the single-lobe kink states $(\phi_{1},\phi_{2})\equiv(\phi_{1,k},\phi_{2,a(k)})\in D_{Z}$ determined by Proposition 4.7 plays a fundamental role. These operators are characterized by the following formal self-adjoint diagonal matrix operator,
$\mathcal{L}_{Z}\mathbb{v}=\Big{(}\Big{(}-c_{j}^{2}\frac{d^{2}}{dx^{2}}v_{j}+\cos(\phi_{j})v_{j}\Big{)}\delta_{j,k}\Big{)},\quad\l 1\leqq j,k\leqq 2,\;\;\mathbb{v}=(v_{j})_{j=1}^{2},\;\;c_{j}>0,$ (1.10)
where $\delta_{j,k}$ denotes the Kronecker symbol, and defined on domains with $\delta$-type interaction at the vertex $\nu=0$, $D(\mathcal{L}_{Z})=D_{Z}$ in (1.7). It is to be observed that the particular family determined by Proposition 4.7 of single-lobe kink states under consideration is such that $(\phi_{1,k},\phi_{2,a(k)})\in D(\mathcal{L}_{Z})$ for every $k$ admissible. An interesting characteristic of the spectrum structure associated with operators in (1.10) on a tadpole graph is that they have a non-trivial Morse index (larger than or equal to 1, in general) which makes the stability study not so immediate. Here we use a novel linear instability criterion for stationary solutions of evolution models on metric graphs (see Theorem 3.2 below) developed by Angulo and Cavalcante [9] (see also [14, 15, 16]) and adapted to the case of the sine-Gordon model on tadpole graphs, as well as a splitting eigenvalue method introduced by Angulo [7] (see Lemma 5.1 below and Angulo [6, 7]). This method is applied to the operator $\mathcal{L}_{Z}\equiv\operatorname{diag}(\mathcal{L}_{1},\mathcal{L}_{a})$ on $D_{Z}$. More precisely, we reduce the eigenvalue problem associated to $\mathcal{L}_{Z}$ into two classes of eigenvalue problems, one for $\mathcal{L}_{1}$ with periodic boundary conditions on $[-L,L]$ and the second one for $\mathcal{L}_{a}$ with $\delta$-type boundary condition on the half-line $[L,\infty)$. Hence, under the restrictions on $L,c_{1},c_{2},Z$ and $k$ given in Theorems 5.10, 5.13 and 5.14, we show that $\ker(\mathcal{L}_{Z})=\{\mathbb{0}\}$ and that the Morse index of $\mathcal{L}_{Z}$ is equal to one. Therefore, as a consequence of the linear instability criterion in Theorem 3.2 we deduce that these single-lobe kink states are linearly unstable by the sine-Gordon flow on a tadpole graph (see Theorem 5.14 below).
Our last main result is that the linear instability of the single-lobe kink states implies also its nonlinear instability (Theorem 5.15 below). The strategy to prove this connection is to use the fact that the mapping data-solution associated to the sine-Gordon model is of class $C^{2}$ on the energy space $\mathcal{E}(\mathcal{G})\times L^{2}(\mathcal{G})$, with
$\mathcal{E}(\mathcal{G})=\{(f,g)\in H^{1}(\mathcal{G}):f(-L)=f(L)=g(L)\},$ (1.11)
$\mathcal{E}(\mathcal{G})$ will be also called the closed continuous subspace at the vertex $\nu=L$ of $H^{1}(\mathcal{G})$. In this way, to carry out this strategy we establish a complete theory on the local well-posedness problem for the sine-Gordon model on $\mathcal{E}(\mathcal{G})\times L^{2}(\mathcal{G})$ in Section 2.
1.1 Main results
Let us summarize the main contributions of this paper:
    First, we prove that the Cauchy problem associated to (1.4) on a tadpole graph $\mathcal{G}$ is well-posed in the energy space $\mathcal{E}(\mathcal{G})\times L^{2}(\mathcal{G})$ with $\mathcal{E}(\mathcal{G})$ defined in (2.1) (see Section 2). This is the content of Theorem 2.1. Even though the well-posedness result is not part of the spectral stability analysis, it is fundamental to reach a nonlinear conclusion; see Theorem 5.15 below.
    In Section 3 we establish a general instability criterion for static solutions for the sine-Gordon model (1.4) on a tadpole graph. The reader can find this result in Theorem 3.2 below. It essentially provides sufficient conditions on the flow of the semigroup generated by the linearization of (1.4) around a priori stationary solutions, for the existence of a pair of positive/negative real eigenvalues of the linearized operator based on its Morse and nullity indexes.
    We provide a complete characterization of the single-lobe kink state profiles $\Theta=(\phi_{1},\phi_{2})$ for the sine-Gordon model on a tadpole graph defined in Definition 1.1 in terms of the parameters $L,c_{1},c_{2}$ and $Z\in(-\infty,\frac{2}{\pi c_{2}})$, in the specific case of the shift-parameter $a<0$ in the kink-profile formula in (1.8). It is shown that $\Theta$ belongs to the domain space $D_{Z}$ in (1.7) only for specific relations between $L,c_{1},c_{2}$, $Z$, and the elliptic modulus $k$ defining profile $\phi_{1,k}$ in (1.9). This implies, in turn, that there does not exist a single-lobe kink state solution with a first-component being of the form in (1.9) (and hence, not satisfying the $\delta$-boundary condition in $D_{Z}$), for a large set of parameters $L,c_{1},c_{2},Z,k$ (see Proposition 4.7).
    We determine (see Section 5) the specific spectral structure for the family of self-adjoint operators $(\mathcal{L}_{Z},D_{Z})$ with $\mathcal{L}_{Z}$ defined in (1.10) and $(\phi_{1},\phi_{2})=(\phi_{1,k},\phi_{2,a(k)})\in D_{Z}$ determined by Proposition 4.7. In particular, we establish a splitting eigenvalue method in Lemma 5.1 which is an essential ingredient to prove that the Morse index of $\mathcal{L}_{Z}$ is exactly one and with a trivial kernel (see Theorem 5.14).
    In Theorem 5.15 we establish that the single-lobe kink state profiles $(\phi_{1,k},\phi_{2,a(k)})\in D_{Z}$ determined by Proposition 4.7, are spectrally and nonlinearly unstable for the sine-Gordon model (1.4) on a tadpole graph. As we mentioned above, the fact that the mapping data-solution associated to the sine-Gordon model on a tadpole graph is of class $C^{2}$ on the energy space $\mathcal{E}(\mathcal{G})\times L^{2}(\mathcal{G})$ is crucial. We also obtained that the degenerate single-lobe kink state $(\pi,\phi_{2,0})\in D_{Z}$ with $Z=\frac{2}{\pi c_{2}}$, $\phi_{2,0}(x)=4\arctan\big{[}e^{-\frac{1}{c_{2}}(x-L)}\big{]}$, $x\in[L,\infty)$, is also nonlinearly unstable.
    For completeness and convenience of the reader, in Appendix A we formulate some tools from the extension theory of symmetric operators by Krein and von Neumann which are suitable for our needs. In particular, we establish a Perron-Frobenius result which is used in the accurate Morse index estimate $n(\mathcal{L}_{Z})\leqq 1$ for $\mathcal{L}_{Z}$ in (1.10) and $(\phi_{1},\phi_{2})$ being a priori single-lobe kink state. Moreover, for completeness in the exposition we also show the general estimate $n(\mathcal{L}_{Z})\leqq 2$ using extension theory.
Notation
[…]
$\quad H^{m}(\mathcal{G})=H^{m}(-L,L)\oplus H^{m}(L,\infty),\quad m\in\mathbb{N},$
with the natural norms. Also, for $\mathbf{u}=(u_{j})_{j=1}^{2}$, $\mathbf{v}=(v_{j})_{j=1}^{2}\in L^{2}(\mathcal{G})$, the inner product is defined by
[…]
2 Well-posedness for the sine-Gordon on a tadpole graph
In this section we show the local well-posedness for the sine-Gordon model (1.4) on a tadpole graph $\mathcal{G}$ in the energy-space $\mathcal{E}(\mathcal{G})\times L^{2}(\mathcal{G})$, with
$\mathcal{E}(\mathcal{G})=\{(f,g)\in H^{1}(\mathcal{G}):f(-L)=f(L)=g(L)\}.$ (2.1)
The space $\mathcal{E}(\mathcal{G})$ represents the closed continuous subspace at the vertex $\nu=L$ of $H^{1}(\mathcal{G})$.
Next, if we denote a wave function $\mathbb{U}$ on the tadpole graph $\mathcal{G}$ by $\mathbb{W}=(\boldsymbol{u},\boldsymbol{v})^{\top}$ with $\boldsymbol{u}=(u_{1},u_{2})^{\top}$, $\boldsymbol{v}=(v_{1},v_{2})^{\top}$, we get the following matrix-form for (1.4),
$\mathbb{W}_{t}=JE\mathbb{W}+F(\mathbb{W})$ (2.2)
with
$J=\left(\begin{array}[]{cc}0&I_{2}\\ -I_{2}&0\end{array}\right),\;E=\left(\begin{array}[]{cc}\mathcal{F}&0\\ 0&I_{2}\end{array}\right),\;F(\mathbb{W})=\left(\begin{array}[]{c}0\\ 0\\ -\sin u_{1}\\ -\sin u_{2}\end{array}\right),$ (2.3)
$I_{2}$ is the identity matrix of order 2 and $\mathcal{F}$ is the diagonal matrix,
$\mathcal{F}=\Big{(}\Big{(}-c_{j}^{2}\frac{d^{2}}{dx^{2}}\Big{)}\delta_{j,k}\Big{)},\quad 1\leqq j,k\leqq 2.$ (2.4)
The main result of this section can be stated as follows.
Theorem 2.1.
The Cauchy problem associated to the sine-Gordon model (2.2) on a tadpole graph $\mathcal{G}$ is locally well-posed on the energy space $\mathcal{E}(\mathcal{G})\times L^{2}(\mathcal{G})$. More precisely, for any $\boldsymbol{\Psi}\in\mathcal{E}(\mathcal{G})\times L^{2}(\mathcal{G})$ there exists $T>0$ such that the sine-Gordon system (2.2) has a unique solution,
$\boldsymbol{W}\in C([0,T];\mathcal{E}(\mathcal{G})\times L^{2}(\mathcal{G})),$
with initial condition $\boldsymbol{W}(0)=\boldsymbol{\Psi}$. Furthermore, for every $T_{0}\in(0,T)$ the data-solution map,
$\boldsymbol{\Psi}\in\mathcal{E}(\mathcal{G})\times L^{2}(\mathcal{G})\mapsto\boldsymbol{W}\in C([0,T_{0}];\mathcal{E}(\mathcal{G})\times L^{2}(\mathcal{G}))$
is, at least, of class $C^{2}$.
The proof of Theorem 2.1 requires some preliminary material. We start by describing the spectrum of the one parameter family of self-adjoint operators $(-\Delta,D_{Z})$ for $Z\in\mathbb{R}$ defined in (1.2) and (1.7) (for a related result in the context of tadpole graphs see Noja and Pelinovsky [42], Proposition A.1). In order to simplify the notation we will write $\mathcal{F}_{Z}$ to denote the operator $-\Delta$ acting on $L^{2}(\mathcal{G})$ with domain $D_{Z}$, for each $Z\in\mathbb{R}$.
Theorem 2.2.
For every $Z\in\mathbb{R}$ the essential spectrum of the self-adjoint operator $\mathcal{F}_{Z}$ is purely absolutely continuous and $\sigma_{\mathrm{\tiny{ess}}}(\mathcal{F}_{Z})=\sigma_{\mathrm{\tiny{ac}}}(\mathcal{F}_{Z})=[0,\infty)$. If $Z>0$ then $\mathcal{F}_{Z}$ has exactly one negative eigenvalue, i.e., its point spectrum is $\sigma_{\mathrm{\tiny{pt}}}(\mathcal{F}_{Z})=\{-\varrho_{Z}^{2}\}$, where $\varrho_{Z}>0$ is the only positive root of the trascendental equation
$\varrho\left(\frac{2}{c_{1}}\tanh\Big{(}\frac{\varrho L}{c_{1}}\Big{)}+\frac{1}{c_{2}}\right)-Z=0,$
and with eigenfunction
$\boldsymbol{\Phi}_{Z}=\begin{pmatrix}\cosh(\varrho_{Z}x/c_{1})\\ e^{-\varrho_{Z}x/c_{2}}\end{pmatrix}.$
If $Z\leqq 0$ then $\mathcal{F}_{Z}$ has no point spectrum, $\sigma_{\mathrm{\tiny{pt}}}(\mathcal{F}_{Z})=\varnothing$.
Proof.
First, it is to be observed that from Theorem A.4 and Proposition A.5 the Schrödinger (symmetric) operator $(-\Delta,D(-\Delta))$ with domain $D(-\Delta)$ (which is defined in (A.5)) and acting on $L^{2}(\mathcal{G})$ has deficiency indices $n_{\pm}(-\Delta)=1$ and consequently the one parameter family of self-adjoint extensions $\mathcal{F}_{Z}$ satisfies $n(\mathcal{F}_{Z})\leqq 1$ for all $Z\in\mathbb{R}$.
Let us now consider the spectral problem $\mathcal{F}_{Z}\mathbf{U}=\lambda\mathbf{U}$ with $\mathbf{U}=(u,v)\in D_{Z}$. In view of the geometry of the tadpole graph $\mathcal{G}$ the spectrum of $\mathcal{F}_{Z}$ is the union of two sets: the set $\lambda$ for which $v\equiv 0$, and the set of $\lambda$ for which $v\neq 0$. In the first case, we have the point spectral problem
[…]
regardless of the value of $Z\in\mathbb{R}$. The eigenvalues of this spectral problem are precisely $\lambda=n^{2}\pi^{2}c_{1}^{2}/L^{2}$ with $n\in\mathbb{N}$, and for each $\lambda$ in this set the eigenfuction of $\mathcal{F}_{Z}$ is given by
$u(x)$ $=\sin\big{(}n\pi x/L\big{)},$ $x\in(-L,L),$
$v(x)$ $=0,$ $x\in(L,\infty),$
for every $Z\in\mathbb{R}$.
The second set includes the absolutely continuous spectrum, $\sigma_{\mathrm{\tiny{ac}}}(\mathcal{F}_{Z})=[0,\infty)$, and for each $\lambda=k^{2}\in[0,\infty)$ the Jost functions of $\mathcal{F}_{Z}$ are given by
$u(x)$ $=a(k)\big{(}e^{ikx/c_{1}}+e^{-ikx/c_{1}}\big{)},$ $x\in(-L,L),$
$v(x)$ $=e^{ikx/c_{2}}+b(k)e^{-ikx/c_{2}},$ $x\in(L,\infty).$
The coefficients $a(k),b(k)$ can be obtained from the boundary conditions in (1.7), yielding the equation
$\mathbf{M}\begin{pmatrix}a(k)\\ b(k)\end{pmatrix}=\mathbf{g}(k),$
where
$\mathbf{M}=\begin{pmatrix}2\cos\big{(}kL/c_{1}\big{)}&-e^{-ikL/c_{2}}\\ \frac{k}{c_{1}}\sin\big{(}kL/c_{1}\big{)}&\big{(}\frac{ik}{c_{2}}-Z\big{)}e^{-ikL/c_{2}}\end{pmatrix},\qquad\mathbf{g}(k)=e^{ikL/c_{2}}\begin{pmatrix}1\\ \big{(}\frac{ik}{c_{2}}+Z\big{)}\end{pmatrix}.$
The determinant of $\mathbf{M}$ is
$|\mathbf{M}|=e^{ikL/c_{2}}\Big{[}2\Big{(}\frac{ik}{c_{2}}-Z\Big{)}\cos\Big{(}\frac{kL}{c_{1}}\Big{)}+\frac{k}{c_{1}}\sin\Big{(}\frac{kL}{c_{1}}\Big{)}\Big{]}.$
A straightforward calculation shows that $|\mathbf{M}|=0$ if and only if $k=0$ and $Z=0$ (yielding in that case the constant solutions $u=v=2a(0)$). Hence for all $k\neq 0$ and all $Z$ the coefficients are given by
$\begin{pmatrix}a(k)\\ b(k)\end{pmatrix}=\Big{[}2\Big{(}\frac{ik}{c_{2}}-Z\Big{)}\cos\Big{(}\frac{kL}{c_{1}}\Big{)}+\frac{k}{c_{1}}\sin\Big{(}\frac{kL}{c_{1}}\Big{)}\Big{]}^{-1}\begin{pmatrix}\frac{2ik}{c_{2}}e^{-ikL/c_{2}}\\ -\frac{k}{c_{1}}\sin\Big{(}\frac{kL}{c_{1}}\Big{)}+2\Big{(}\frac{ik}{c_{2}}+Z\Big{)}\cos\Big{(}\frac{kL}{c_{1}}\Big{)}\end{pmatrix}.$
Since $a(k)$ and $b(k)$ are bounded and non-zero, there are no spectral singularities in the absolutely continuous spectrum of $\mathcal{F}_{Z}$ in $L^{2}(\mathcal{G})$.
Finally, let us examine whether this second set includes isolated negative eigenvalues, $\lambda<0$, with $v\neq 0$. In such a case, since $v\in H^{2}([L,\infty))$ implies $v\to 0$ as $x\to\infty$, we consider an eigenfunction of the form
$u(x)$ $=\cosh(\sqrt{|\lambda|}x/c_{1}),$ $x\in(-L,L),$
$v(x)$ $=e^{-\sqrt{|\lambda|}x/c_{2}},$ $x\in(L,\infty),$
under the assumption $\lambda<0$. The boundary conditions in (1.7) yield the system
$\cosh(\sqrt{|\lambda|}L/c_{1})$ $=e^{-\sqrt{|\lambda|}L/c_{2}},$
$\frac{2\sqrt{|\lambda|}}{c_{1}}\sinh(\sqrt{|\lambda|}L/c_{1})$ $=\Big{(}Z-\frac{\sqrt{|\lambda|}}{c_{2}}\Big{)}e^{-\sqrt{|\lambda|}L/c_{2}}.$
Upon substitution we obtain the trascendental equation
$\sqrt{|\lambda|}\left(\frac{2}{c_{1}}\tanh\Big{(}\frac{\sqrt{|\lambda|}L}{c_{1}}\Big{)}+\frac{1}{c_{2}}\right)=Z,$
which has a unique real positive solution $\sqrt{|\lambda|}>0$ only in the case when $Z>0$. Indeed, consider the function
$G(\varrho):=\varrho\left(\frac{2}{c_{1}}\tanh\Big{(}\frac{\varrho L}{c_{1}}\Big{)}+\frac{1}{c_{2}}\right)-Z,\qquad\varrho\in[0,\infty),$
assuming $Z>0$. Clearly, $G(0)=-Z<0$ and $G(\varrho)\to\infty$ as $\varrho\to\infty$. Moreover,
$G^{\prime}(\varrho)=\frac{2}{c_{1}}\tanh\Big{(}\frac{\varrho L}{c_{1}}\Big{)}+\frac{1}{c_{2}}+\frac{2\varrho L}{c_{1}^{2}}\frac{1}{\cosh^{2}(\varrho L/c_{1})}>0,$
for all $\varrho\in(0,\infty)$ and, thus, $G$ is strictly increasing. By the intermediate value theorem there exists a unique value $\varrho_{Z}>0$ such that $G(\varrho_{Z})=0$.
From the previous observation we conclude that if $Z>0$ then $n(\mathcal{F}_{Z})=1$, whereas if $Z\leqq 0$ then $n(\mathcal{F}_{Z})=0$. The theorem is proved. ∎
Let us now characterize the resolvent of the operator $\mathcal{A}=JE$ (defined in (2.3)) in $H^{1}(\mathcal{G})\times L^{2}(\mathcal{G})$. For that purpose we use the description of the spectrum of $\mathcal{F}_{Z}$ from Theorem 2.2.
Theorem 2.3.
Let $Z\in\mathbb{R}$. For any $\lambda\in\mathbb{C}$ such that $-\lambda^{2}\in\rho(\mathcal{F}_{Z})$ we have that $\lambda$ belongs to the resolvent set of $\mathcal{A}=JE$, with $D(\mathcal{A})=D_{Z}\times L^{2}(\mathcal{G})$; moreover, the resolvent,
$R(\lambda:\mathcal{A})=(\lambda I_{4}-\mathcal{A})^{-1}\,:\,H^{1}(\mathcal{G})\times L^{2}(\mathcal{G})\to D(\mathcal{A}),$
has the following representation for $\mathbf{U}=(\boldsymbol{u},\boldsymbol{v})$,
$R(\lambda:\mathcal{A})\mathbf{U}=\begin{pmatrix}-R(-\lambda^{2}:\mathcal{F}_{Z})(\boldsymbol{v}+\lambda\boldsymbol{u})\\ -\lambda R(-\lambda^{2}:\mathcal{F}_{Z})(\boldsymbol{v}+\lambda\boldsymbol{u})-\boldsymbol{u}\end{pmatrix},$ (2.5)
where $R(-\lambda^{2}:\mathcal{F}_{Z})=(-\lambda^{2}I_{2}-\mathcal{F}_{Z})^{-1}:L^{2}(\mathcal{G})\to D_{Z}$.
Proof.
The proof is very similar to that of Theorem 2.2 in [14], so we gloss over some of the details. Suppose that for $\mathbf{U}=(\boldsymbol{u},\boldsymbol{v})\in H^{1}(\mathcal{G})\times L^{2}(\mathcal{G})$ there holds the relation $(\lambda I_{4}-\mathcal{A})^{-1}=(\boldsymbol{\phi},\boldsymbol{\psi})\in D(\mathcal{A})$. This relation holds if and only if
$\boldsymbol{u}$ $=\lambda\boldsymbol{\phi}-\boldsymbol{\psi},$
$\boldsymbol{v}$ $=\mathcal{F}_{Z}\boldsymbol{\phi}+\lambda\boldsymbol{\psi}.$
Upon substitution we obtain $(\mathcal{F}_{Z}+\lambda^{2}I_{2})\boldsymbol{\phi}=\boldsymbol{v}+\lambda\boldsymbol{u}$. But, by hypothesis, $-\lambda^{2}\in\rho(\mathcal{F}_{Z})$ so that $\boldsymbol{\phi}=-R(-\lambda^{2}:\mathcal{F}_{Z})(\boldsymbol{v}+\lambda\boldsymbol{u})$ and, consequently, $\boldsymbol{\psi}=-\lambda R(-\lambda^{2}:\mathcal{F}_{Z})(\boldsymbol{v}+\lambda\boldsymbol{u})-\boldsymbol{u}$, as claimed. ∎
It is actually possible to provide some explicit formulae for $R(-\lambda^{2}:\mathcal{F}_{Z})$ in the case where $-\lambda^{2}\in\rho(\mathcal{F}_{Z})$ and $\lambda\in\mathbb{R}$.
Proposition 2.4.
Fix $Z\in\mathbb{R}$ and consider $-\lambda^{2}\in\rho(\mathcal{F}_{Z})$ with $\lambda\in\mathbb{R}$. For any $\boldsymbol{u}=(u_{1},u_{2})\in L^{2}(\mathcal{G})$ we have:
  • (a)
    In the case when $Z\leq 0$ and for all $\lambda>0$ (without loss of generality) such that $-\lambda^{2}\in\rho(\mathcal{F}_{Z})$, set $\Phi:=(\mathcal{F}_{Z}+\lambda^{2}I_{2})^{-1}\boldsymbol{u}$ to obtain
[…]
  • (b)
    In the case when $Z>0$ and for all $\lambda>0$ (without loss of generality) such that $-\lambda^{2}\in\rho(\mathcal{F}_{Z})$ and with $-\lambda\neq\varrho_{Z}^{2}$, set $\Psi:=(\mathcal{F}_{Z}+\lambda^{2}I_{2})^{-1}\boldsymbol{u}$ to obtain
    $\Psi_{1}(x)$ $=\frac{\cosh(\varrho_{z}x/c_{1})}{\lambda+\varrho_{z}^{2}}\big{\langle}u_{1},\cosh(\varrho_{z}x/c_{1})\big{\rangle}+\Phi_{1}(x),$ $x\in(-L,L),$ (2.7)
    $\Psi_{2}(x)$ $=\frac{e^{-\varrho_{z}x/c_{2}}}{\lambda+\varrho_{z}^{2}}\big{\langle}u_{2},e^{-\varrho_{z}x/c_{2}}\big{\rangle}+\Phi_{2}(x),$ $x\in(L,\infty).$
The constants $d_{j}=d_{j}(\lambda,\Phi_{j})$, $j=1,2$, are chosen such that $\Phi\in D_{Z}$.
Proof.
Formulae (2.6) and (2.7) result from the variation of constants formula: a solution to the homogeneous problem plus a convolution with the Green’s function associated to the operator (see formula (2.6) when $Z\leqq 0$). In the case when there exists a negative eigenvalue (that is, when $Z>0$) one also needs to project onto the associated eigenspace (formula (2.7)). The reader can directly verify that $(\mathcal{F}_{Z}+\lambda^{2}I_{2})\Phi=\boldsymbol{u}$ if $Z\leqq 0$, and that $(\mathcal{F}_{Z}+\lambda^{2}I_{2})\Psi=\boldsymbol{u}$ if $Z>0$, whenever $-\lambda^{2}\in\rho(\mathcal{F}_{Z})$. We have specialized both formulae for $\lambda>0$ for convenience, without loss of generality. These formulae are particular cases of the well-known Krein’s formula for the resolvents of self-adjoint extensions of a symmetric operator (cf. Albeverio et al. [4], Appendix A; Albeverio and Kurasov [5], §1.2). For the Laplacian case see, for example, Theorem 3.1.2 in Albeverio et al. [4], p. 77, or Angulo and Ponce [17], eq. (3.13), for the periodic case. ∎
Next, from Theorem 2.2 we can define the following equivalent $X^{1}_{Z}$-norm to $H^{1}(\mathcal{G})$, for $\boldsymbol{v}=(v_{j})_{j=1}^{2}\in H^{1}(\mathcal{G})$
$\|\boldsymbol{v}\|_{X^{1}_{Z}}^{2}=\|\boldsymbol{v}^{\prime}\|^{2}_{L^{2}(\mathcal{G})}+(\beta+1)\|\boldsymbol{v}\|^{2}_{L^{2}(\mathcal{G})}+Z|v_{1}(L)|^{2},$ (2.8)
where $\beta\geqq 0$ is defined as,
$\beta:=\begin{cases}\varrho_{Z}^{2},&\text{if }Z>0,\\ 0,&\text{if }Z\leqq 0.\end{cases}$ (2.9)
We will denote by $H^{1}_{Z}(\mathcal{G})$ the space $H^{1}(\mathcal{G})$ with the norm $\|\cdot\|_{X^{1}_{Z}}$. Moreover, the following well-defined inner product in $H^{1}_{Z}(\mathcal{G})$,
[…]
induces the $X^{1}_{Z}$-norm above (here we are considering $c_{j}^{2}=1$, $j=1,2$, without loss of generality).
The following theorem verifies that the operator $\mathcal{A}=JE$ is indeed the infinitesimal generator of a $C_{0}$-semigroup upon application of the classical Lumer-Phillips theory.
Theorem 2.5.
Let $Z\in\mathbb{R}$ and consider the linear operators $J$ and $E$ defined in (2.3). Then, $\mathcal{A}=JE$ with $D(\mathcal{A})=D(\mathcal{F}_{Z})\times\mathcal{E}(\mathcal{\mathcal{G}})$ is the infinitesimal generator of a $C_{0}$-semigroup $\{\mathcal{S}(t)\}_{t\geqq 0}$ on $H^{1}(\mathcal{\mathcal{G}})\times L^{2}(\mathcal{\mathcal{G}})$. The initial value problem
$\begin{cases}\boldsymbol{w}_{t}=\mathcal{A}\boldsymbol{w},\\ \boldsymbol{w}(0)=\boldsymbol{w}_{0}\in D(\mathcal{A})=D(\mathcal{F}_{Z})\times\mathcal{E}(\mathcal{\mathcal{G}}),\end{cases}$ (2.11)
has a unique solution $\boldsymbol{w}\in C([0,\infty):D(\mathcal{A}))\cap C^{1}([0,\infty):H^{1}(\mathcal{\mathcal{G}})\times L^{2}(\mathcal{\mathcal{G}}))$ given by $\boldsymbol{w}(t)=\mathcal{S}(t)\boldsymbol{w}_{0}$, $t\geqq 0$. Moreover, for any $\boldsymbol{u}\in H^{1}(\mathcal{\mathcal{G}})\times L^{2}(\mathcal{\mathcal{G}})$ and $\theta>\beta+1$, where $\beta\geqq 0$ is given in (2.9), we have the representation formula
$\mathcal{S}(t)\boldsymbol{u}=\frac{1}{2\pi i}\int_{\theta-i\infty}^{\theta+i\infty}e^{\lambda t}R(\lambda:\mathcal{A})\boldsymbol{u}\,d\lambda$ (2.12)
where $\lambda\in\rho(\mathcal{A})$ with $\mathrm{Re}\,\lambda=\theta$ and $R(\lambda:\mathcal{A})=(\lambda I-\mathcal{A})^{-1}$, and for every $\delta>0$, the integral converges uniformly in $t$ for every $t\in[\delta,1/\delta]$.
Proof.
For convenience of the reader we sketch the main steps of the proof (without lost of generality we consider $c_{j}^{2}=1$):
  1. (1)
    We consider the Hilbert space $X_{Z}\equiv H^{1}_{Z}(\mathcal{\mathcal{G}})\times L^{2}(\mathcal{\mathcal{G}})$ with inner product $\langle\cdot,\cdot\rangle_{X_{Z}}=\langle\cdot,\cdot\rangle_{1,Z}+\langle\cdot,\cdot\rangle$, with $\langle\cdot,\cdot\rangle_{1,Z}$ defined in (2.10). Define $\mathcal{B}=\mathcal{A}-\gamma I$, $\gamma=\beta+1>0$, with $\beta\geqq 0$ as in (2.9). Then the following linear initial value problem
    $\begin{cases}\boldsymbol{u}_{t}=\mathcal{B}\boldsymbol{u}\\ \boldsymbol{u}(0)=\boldsymbol{u}_{0}\in D(\mathcal{B})=D(\mathcal{F}_{Z})\times\mathcal{E}(\mathcal{\mathcal{G}})\end{cases}$ (2.13)
    has a unique solution $\boldsymbol{u}\in C([0,\infty):D(\mathcal{B}))\cap C^{1}([0,\infty):X_{Z})$ given by $\boldsymbol{u}(t)=\mathcal{U}(t)\boldsymbol{u}_{0}$, $t\geqq 0$, where $\{\mathcal{U}(t)\}_{t\geqq 0}$ is a $C_{0}$-semigroup of contractions on $X_{Z}$. Indeed, first we observe that $\mathcal{B}$ is dissipative on $X_{Z}$: for $\mathbf{U}=(\boldsymbol{u},\boldsymbol{v})\in D(\mathcal{B})$ with $\boldsymbol{u}=(u_{j})_{j=1}^{2}\in D(\mathcal{F}_{Z})$ and $\boldsymbol{v}=(v_{j})_{j=1}^{2}\in\mathcal{E}(\mathcal{\mathcal{G}})$ (with $u_{j},v_{j}$ real-valued without lost of generality)
    $\langle-\mathcal{B}\mathbf{U},\mathbf{U}\rangle_{X_{Z}}=(\beta+1)\Big{[}\|\boldsymbol{u}\|^{2}_{X^{1}_{Z}}+\|\boldsymbol{v}\|^{2}_{L^{2}(\mathcal{\mathcal{G}})}-\langle\boldsymbol{u},\boldsymbol{v}\rangle\Big{]}\geqq 0,$ (2.14)
    because of the Cauchy-Schwartz inequality and $\|\boldsymbol{u}\|^{2}_{L^{2}(\mathcal{\mathcal{G}})}\leqq\|\boldsymbol{u}\|^{2}_{X^{1}_{Z}}$ by (2.8). Next, we notice that from Theorem 2.3 we can choose $\lambda$ such that $\lambda+\gamma>0$ and $\lambda+\gamma\in\rho(\mathcal{A})$. Hence the range, $R(\lambda I-\mathcal{B})=R((\lambda+\gamma)I-\mathcal{A})$, of $\lambda I-\mathcal{B}$ is $X_{Z}$. Henceforth, upon application of the classical Lumer-Phillips Theorem (see, e.g., Pazy [43]) we conclude that $\mathcal{B}$ is the infinitesimal generator of a $C_{0}$-semigroup of contractions $\{\mathcal{U}(t)\}_{t\geqq 0}$ on $X_{Z}$, as claimed.
  2. (2)
    Now define $\mathcal{S}(t)=e^{\gamma t}\mathcal{U}(t)$. Then $\{\mathcal{S}(t)\}_{t\geqq 0}$ is a $C_{0}$-semigroup on $X_{Z}$ with infinitesimal generator $\mathcal{S}^{\prime}(0)=\gamma I+\mathcal{A}-\gamma I=\mathcal{A}$. Then, since the norm $\|\cdot\|_{H^{1}(\mathcal{\mathcal{G}})}$ is equivalent to the norm $\|\cdot\|_{X_{Z}^{1}}$ on $H^{1}(\mathcal{\mathcal{G}})$, we obtain that $\{\mathcal{S}(t)\}_{t\geqq 0}$ is a $C_{0}$-semigroup on $H^{1}(\mathcal{\mathcal{G}})\times L^{2}(\mathcal{\mathcal{G}})$ and $\boldsymbol{w}(t)=\mathcal{S}(t)\boldsymbol{w}_{0}$ is the unique solution for the linear problem (2.11). Moreover, $\|\mathcal{S}(t)\|_{H^{1}(\mathcal{\mathcal{G}})\times L^{2}(\mathcal{\mathcal{G}})}\leqq Me^{\gamma t}$, for $M>0$, $\gamma=\beta+1>0$ and $t\geqq 0$. Therefore, from Theorem 2.3, standard semigroup theory and properties of the Laplace transform, we obtain for $\theta>\beta+1$ the representation formula in (2.12). This finishes the proof.
Next result simply states the (expected) invariance property of the energy space under the action of the semigroup.
Proposition 2.6.
The subspace $\mathcal{E}(\mathcal{G})\times L^{2}(\mathcal{G})$ is invariant under the semigroup $\{\mathcal{S}(t)\}_{t\geqq 0}$ defined by formula (2.12). Moreover, $\mathcal{S}(t)(\mathcal{E}(\mathcal{G})\times L^{2}(\mathcal{G}))\subset\mathcal{E}(\mathcal{G})\times\mathcal{C}(\mathcal{G})$, $t>0$, where
$\mathcal{C}(\mathcal{G})=\{(v_{j})_{j=1}^{2}\in L^{2}(\mathcal{G}):v_{1}(L)=v_{2}(L)\}.$ (2.15)
Proof.
By the representation of $\mathcal{S}(t)$ in (2.12) it suffices to show that the resolvent operator $R(\lambda:\mathcal{A})\Phi\in\mathcal{E}(\mathcal{G})\times L^{2}(\mathcal{G})$ for $\Phi\in\mathcal{E}(\mathcal{G})\times L^{2}(\mathcal{G})$. Indeed, for $\Psi=(\mathbb{u},\mathbb{v})$ we have from formula (2.5) that $R(-\lambda^{2}:\mathcal{F}_{Z})(\lambda\mathbb{u}+\mathbb{v})\in D(\mathcal{F}_{Z})\subset\mathcal{E}(\mathcal{G})$ and so $R(\lambda:\mathcal{A})\Phi\in\mathcal{E}(\mathcal{G})\times\mathcal{E}(\mathcal{G})\subset\mathcal{E}(\mathcal{G})\times\mathcal{C}(\mathcal{G})\subset\mathcal{E}(\mathcal{G})\times L^{2}(\mathcal{G})$. ∎
We are now ready to prove Theorem 2.1.
Proof of Theorem 2.1.
Based on Theorem 2.5 and Proposition 2.6, the local well-posedness result in $\mathcal{E}(\mathcal{G})\times L^{2}(\mathcal{G})$ follows from the Banach fixed point theorem and standard arguments. It is to be noticed that the contraction mapping principle has the advantage that, since the nonlinear term $F(z)$ is smooth, then the Implicit Function Theorem implies the smoothness of the data-solution map associated to the sine-Gordon equation (see [11, 14, 15, 16] for further information). Lastly, from Proposition 2.6 we obtain that for all $t\in(0,T]$, $\boldsymbol{w}(t)\in\mathcal{E}(\mathcal{G})\times\mathcal{C}(\mathcal{Y})$. This finishes the proof. ∎
3 Linear instability criterion for the sine-Gordon model on a tadpole graph
In this section we establish a linear instability criterion of stationary solutions for the sine-Gordon model (2.2) on a tadpole graph $\mathcal{G}$. More importantly, the criterion also applies to any type of stationary solutions independently of the boundary conditions under consideration and can be therefore used to study configurations with boundary rules at the vertex of $\delta^{\prime}$-interaction type, or with other types of stationary solutions to the sine-Gordon equation.
Let us suppose that $JE$ on a domain $D(JE)\subset L^{2}(\mathcal{G})$ is the infinitesimal generator of a $C_{0}$-semigroup on $L^{2}(\mathcal{G})$ and the stationary solution $\Phi=(\phi_{1}(x),\phi_{2}(x),0,0)\in D(JE)$. Thus, every component satisfies the equation
$-c^{2}_{j}\phi^{\prime\prime}_{j}+\sin\phi_{j}=0,\quad j=1,2.$ (3.1)
Now, we suppose that $\mathbb{W}$ satisfies formally equality in (2.2) and we define
$\mathbb{V}\equiv\mathbb{W}-\Phi.$ (3.2)
Then, from (3.1) we obtain the following linearized system for (2.2) around $\Phi$,
$\mathbb{V}_{t}=J\mathcal{E}\mathbb{V},$ (3.3)
with $\mathcal{E}$ being the $4\times 4$ diagonal-matrix $\mathcal{E}=\left(\begin{array}[]{cc}\mathcal{L}&0\\ 0&I_{2}\end{array}\right)$, and
$\mathcal{L}=\Big{(}\Big{(}-c_{j}^{2}\frac{d^{2}}{dx^{2}}+\cos\phi_{j}\Big{)}\delta_{j,k}\Big{)},\qquad 1\leqq j,k\leqq 2.$ (3.4)
We point out the equality $J\mathcal{E}=JE+\mathcal{T}$, with
$\mathcal{T}=\left(\begin{array}[]{cc}0&0\\ \big{(}-\cos(\phi_{j})\,\delta_{j,k}\big{)}&0\end{array}\right)$
being a bounded operator on $H^{1}(\mathcal{G})\times L^{2}(\mathcal{G})$. This implies that $J\mathcal{E}$ also generates a $C_{0}$-semigroup on $H^{1}(\mathcal{G})\times L^{2}(\mathcal{G})$ (see Pazy [43]).
In the sequel, our objective is to provide sufficient conditions for the trivial solution $\mathbb{V}\equiv 0$ to be unstable by the linear flow (3.3). More precisely, we are interested in finding a growing mode solution of (3.3) with the form $\mathbb{V}=e^{\lambda t}\Psi$ and $\mathrm{Re}\,\lambda>0$. In other words, we need to solve the formal system
$J\mathcal{E}\Psi=\lambda\Psi,$ (3.5)
with $\Psi\in D(J\mathcal{E})$. If we denote by $\sigma(J\mathcal{E})=\sigma_{\mathrm{pt}}(J\mathcal{E})\cup\sigma_{\mathrm{ess}}(J\mathcal{E})$ the spectrum of $J\mathcal{E}$ (namely, $\lambda\in\sigma_{\mathrm{pt}}(J\mathcal{E})$ if $\lambda$ is isolated and with finite multiplicity), the later discussion suggests the usefulness of the following definition.
Definition 3.1.
The stationary vector solution $\Phi\in D(\mathcal{E})$ is said to be spectrally stable for model sine-Gordon if the spectrum of $J\mathcal{E}$, $\sigma(J\mathcal{\mathcal{missing}}E)$, satisfies $\sigma(J\mathcal{E})\subset i\mathbb{R}.$ Otherwise, the stationary solution $\Phi\in D(\mathcal{E})$ is said to be spectrally unstable.
It is standard to show that $\sigma_{\mathrm{pt}}(J\mathcal{E})$ is symmetric with respect to both the real and imaginary axes and that $\sigma_{\mathrm{ess}}(J\mathcal{E})\subset i\mathbb{R}$, by supposing $J$ skew-symmetric and $\mathcal{E}$ self-adjoint (for instance, under Assumption $(S_{3})$ below for $\mathcal{L}$; see [28, Lemma 5.6 and Theorem 5.8]). These cases on $J$ and $\mathcal{E}$ will be considered in our theory. Hence it is equivalent to say that $\Phi\in D(J\mathcal{E})$ is spectrally stable if $\sigma_{\mathrm{pt}}(J\mathcal{E})\subset i\mathbb{R}$, and it is spectrally unstable if $\sigma_{\mathrm{pt}}(J\mathcal{E})$ contains a point $\lambda$ with $\mathrm{Re}\,\lambda>0$.
From (3.5), our eigenvalue problem to solve is now reduced to,
$J\mathcal{E}\Psi=\lambda\Psi,\quad\mathrm{Re}\,\lambda>0,\;\;\Psi\in D(\mathcal{E}).$ (3.6)
Next, we establish our theoretical framework and the main assumptions in order to obtain a nontrivial solution to problem (3.6):
  1. ($S_{1}$)
    $J\mathcal{E}$ is the generator of a $C_{0}$-semigroup $\{S(t)\}_{t\geqq 0}$.
  2. ($S_{2}$)
    Let $\mathcal{L}$ be the matrix-operator in (3.4) defined on a domain $D(\mathcal{L})\subset L^{2}(\mathcal{G})$ on which $\mathcal{L}$ is self-adjoint.
  3. ($S_{3}$)
    Suppose $\mathcal{L}:D(\mathcal{L})\to L^{2}(\mathcal{G})$ is invertible with Morse index $n(\mathcal{L})=1$ and such that $\sigma(\mathcal{L})=\{\lambda_{0}\}\cup J_{0}$ with $J_{0}\subset[r_{0},\infty)$, for $r_{0}>0$ and $\lambda_{0}<0$.
Our linear instability criterion is the following.
Theorem 3.2.
Suppose the assumptions $(S_{1})$ $(S_{3})$ hold. Then the operator $J\mathcal{E}$ has a real positive and a real negative eigenvalue.
The proof of Theorem 3.2 is based in ideas from Lopes [35], Angulo and Cavalcante [8, 9] (see also Angulo and Plaza [14] for a complete proof of this), applied to the case of a tadpole graph.
It is widely known that the spectral instability of a specific traveling wave solution of an evolution type model is a key prerequisite to show their nonlinear instability property (see [28] and references therein). Thus we have the following definition.
Definition 3.3.
A stationary vector solution $\Phi\in D(\mathcal{E})$ is said to be nonlinearly unstable in $X\equiv H^{1}(\mathcal{G})\times L^{2}(\mathcal{G})$-norm for model sine-Gordon if there is $\epsilon>0$, such that for every $\delta>0$ there exist an initial datum $\mathbb{w}_{0}$ with $\|\Phi-\mathbb{w}_{0}\|_{X}<\delta$ and an instant $t_{0}=t_{0}(\mathbb{w}_{0})$ such that $\|\mathbb{w}(t_{0})-\Phi\|_{X}>\epsilon$, where $\mathbb{w}=\mathbb{w}(t)$ is the solution of the sine-Gordon model with $\mathbb{w}(0)=\mathbb{w}_{0}$.
4 Existence of single-lobe kink state for the sine-Gordon equation on a tadpole graph
In this section we show the existence of static solutions of single-lobe kink type for the sine-Gordon model determined by a $\delta$-interaction type at the vertex $\nu=L$ of a tadpole graph (see (1.7)). We will find that these profiles exist as long as we have several restrictions involved the a priori parameters $c_{i}$ and $Z$. We start with the strength value $Z$.
4.1 A priori restrictions on the parameter $Z$
We consider a single-lobe kink state $(\phi_{1},\phi_{2})$ for (1.6). Then, as $\phi_{1}(L)=\phi_{2}(L)>0$ follows that $\phi_{1}$ satisfies the following condition,
$\phi^{\prime}_{1}(L)-\phi^{\prime}_{1}(-L)=\Big{[}\frac{\phi^{\prime}_{2}(L)}{\phi_{2}(L)}+Z\Big{]}\phi_{1}(L)\equiv\alpha_{{}_{Z}}\phi_{1}(L).$ (4.1)
Next, from the fact that $\phi_{1}$ is even we have $2\phi_{1}^{\prime}(L)=\alpha_{Z}\phi_{1}(L)$, and from $\phi^{\prime}_{1}(L)<0$ we get $\alpha_{Z}<0$. Now, from (1.8) we obtain for $y=e^{-a/c_{2}}$
$-Z>\frac{\phi^{\prime}_{2}(L)}{\phi_{2}(L)}=-\frac{1}{c_{2}}\frac{y}{(1+y^{2})\arctan(y)}\equiv f(y),\;\;\quad y>0.$ (4.2)
Thus, we have the following proposition.
Proposition 4.1.
We consider a single-lobe kink state $(\phi_{1},\phi_{2})\in D_{Z}$ for (1.6). Then, the strength value $Z$ satisfies $Z\in(-\infty,\frac{2}{\pi c_{2}})$.
Proof.
From (4.2) we get initially that $Z<\frac{1}{c_{2}}$ because $f(y)>-\frac{1}{c_{2}}$ for all $y>0$. Now, for $a\leqq 0$ we have $y=e^{-a/c_{2}}\geqq 1$. Then $-Z>f(y)\geqq f(1)=-\frac{2}{\pi c_{2}}$, and so $Z<\frac{2}{\pi c_{2}}$. For $a>0$, we have $y<1$ and so $f(y)<f(1)$. Thus, for $a\to 0^{+}$, $f(y)\to f(1)$, and therefore $f(1)\leqq-Z$. But, for $a=0$, from (4.2) follows $-Z>f(1)$. In conclusion, we need to have the restriction $Z\in(-\infty,\frac{2}{\pi c_{2}})$. ∎
4.2 Existence of positive single-lobe states (subluminal librations)
In the following we show the existence of positive solutions for the equation
$-c_{1}^{2}\phi^{\prime\prime}_{1}(x)+\sin(\phi_{1}(x))=0,\qquad x\in[-L,L],$ (4.3)
with a single-lobe on $[-L,L]$, namely, $\phi_{1}$ is even, the maximum of $\phi_{1}$ is achieved at a single internal point of $[-L,L]$, namely, $x=0$, and $\phi^{\prime}(x)<0$ on $(0,L]$ (see Figure 3 and the phase-plane in Figure 5).
Figure 5: Phase-plane of a positive single-lobe kink state for (1.6). Here we take $c_{1}=c_{2}=1$ and the translation variable $\xi=x-ct$ with $c=0$ (subluminal waves). The picture shows the phase plane of the solutions $\phi=\phi(\xi)$ for (1.6). The separatrices are represented in red solid lines and the librations are the periodic waves inside the separatrices (in light gray). The single-lobe kink profile (depicted with blue thick arrow lines) is composed of a periodic libration, symmetric in $[-L,L]$ and connecting $\phi=\pi$ with itself, glued with a decaying kink connecting $\phi=\pi$ to $\phi=0$ along a separatrix (color online).
The solutions to (4.3) of our interest are called subluminal librations, namely, the profiles inside the separatrix induced by the kink-profile $\phi_{2}$ in (1.8) (see Figure 5) which are associated to a traveling speed satisfying $c^{2}<c_{j}^{2}$, where $c_{j}$ is the corresponding characteristic speed on each edge $E_{j}$, which are fixed (see, e.g., Jones et al. [31] for a complete description of subluminal librations). These profiles constitute one of the components of the periodic stationary profiles determined by the sine-Gordon model. Then, by considering the quadratic form associated to (4.3) and by integrating by parts once we get from condition $\phi_{1}^{\prime}(0)=0$,
$-\frac{c_{1}^{2}}{2}[\phi^{\prime}_{1}(x)]^{2}-\cos(\phi_{1}(x))=-\cos(\phi_{1}(0))\Longrightarrow-\frac{c_{1}^{2}}{2}[\phi^{\prime}_{1}(x)]^{2}+1-\cos(\phi_{1}(x))=E,\;\;x\in[-L,L],$ (4.4)
with $E\equiv 1-\cos(\phi_{1}(0))$ is so-called energy-level. Thus, being $\phi_{1}$ a single-lobe profile we need to have $\phi_{1}(0)\in(\pi,2\pi)$ and therefore $E\in(0,2)$.
Next, we consider the following change of variables of snoidal-type (see Figure 7),
$\cos(\phi_{1}(x))=-1+\beta\operatorname{sn}^{2}(\lambda(x-x_{0});k),$ (4.5)
where $\text{sn}(\cdot;k)$ represents the Jacobian elliptic function snoidal determined by the parameter $k\in(0,1)$ (called the elliptic modulus) and $x_{0},\lambda,\beta\in\mathbb{R}$ to be determined. Thus, it is not difficult to see from (4.4) that we arrive to the following equality for all $x\in[-L,L]$,
$2(E-2)\beta\operatorname{sn}^{2}+\beta^{2}[2\beta-(E-2)]\operatorname{sn}^{4}-\beta^{3}\operatorname{sn}^{6}=-2c_{1}^{2}\beta^{2}\lambda^{2}\operatorname{sn}^{2}+2\beta^{2}\lambda^{2}c_{1}^{2}(1+k^{2})\operatorname{sn}^{4}-2\beta^{2}\lambda^{2}c_{1}^{2}k^{2}\operatorname{sn}^{6},$
and therefore $\beta=2-E>0$, $\lambda=\frac{1}{c_{1}}$, and $k=\sqrt{\frac{2-E}{2}}$.
Now, since $\cos(\phi_{1}(0))=1-E$ (see (4.4)) we get from (4.5) that $(2-E)\operatorname{sn}^{2}(-\frac{x_{0}}{c_{1}};k)=2-E$ and therefore
$\operatorname{sn}^{2}(-\frac{x_{0}}{c_{1}};k)=1\;\;\text{implies}\;\;x_{0}=\pm c_{1}(2n+1)K(k),\;n\geqq 0,$
where $K(k)$ is the complete elliptic integral of the first kind defined by
[…]
with $K^{\prime}(k)>0$, $K(k)\to\frac{\pi}{2}$ as $k\to 0$ and $K(k)\to\infty$ as $k\to 1$ (see [22]). We recall that $\operatorname{sn}(\cdot;k)$ is periodic-odd of period $4K(k)$ and, thus, $\operatorname{sn}^{2}(\cdot;k)$ is periodic-even of period $2K(k)$. Hence we can choose any sign for defining $x_{0}$ and $x_{0}=c_{1}K(k)$. Therefore, we get the formula for $\phi_{1}$ as
$\cos(\phi_{1}(x))=-1+2k^{2}\operatorname{sn}^{2}\Big{(}\frac{x}{c_{1}}+K(k);k\Big{)},\quad k\in(0,1).$ (4.6)
We note immediately that $\Phi(x)=\arccos\Big{[}-1+2k^{2}\operatorname{sn}^{2}\Big{(}\frac{x}{c_{1}}+K(k);k\Big{)}\Big{]}$ does not represent the first component of a single-lobe kink profile according to Definition 1.1 (see Figures 6 and 3). The stability properties of these $\Phi$-profiles and of the anti-kink profiles for the sine-Gordon equation will be studied in a future work.
Figure 6: Profile for $\Phi=\Phi(x)$ with $x\in[-\frac{\pi}{4},\frac{\pi}{4}]$, $c_{1}=1$, and any $k\in(0,1)$.
Figure 7: Graphs of the elliptic functions $\operatorname{sn}(x,k)$, $\mathrm{cn}(x,k)$ and $\mathrm{dn}(x,k)$ for $x\in[0,4K(k)]$, $k=0.4$.
Before establishing our existence result of the single-lobe profile according to Definition 1.1, we obtain some relations between the modulus $k^{2}=\frac{2-E}{2}$ and the maximum value of a priori single-lobe $\phi_{1}$ (we recall $\phi_{1}(0)\in(\pi,2\pi)$; see Figure 5):
  1. (i)
    Let $k$ be fixed and $k^{2}\leqq\frac{1}{2}$: $E\in[1,2)$, $\cos(\phi_{1}(0))=-1+2k^{2}\in(-1,0]$ if and only if $\phi_{1}(0)\in(\pi,\frac{3\pi}{2}]$.
  2. (ii)
    For $k$ fixed and $k^{2}>\frac{1}{2}$: $E\in(0,1)$, $\cos(\phi_{1}(0))=-1+2k^{2}\in(0,1)$ if and only if $\phi_{1}(0)\in(\frac{3\pi}{2},2\pi]$.
In the sequel we obtain two explicit formulae for the profile $\phi_{1}$ in (4.6) being an even single-lobe profile on $[-L,L]$ such that $\phi_{1}^{\prime}(x)<0$ for $x\in(0,L)$ and $\phi_{1}(-L)=\phi_{1}(L)=\phi_{2,a}(L)$. As we shall see, these formulae will depend on the position of $\phi_{2,a}(L)$ with regard to $\pi$, where the shift-parameter $a$ determines the kink-soliton profile in (1.8) (see Figure 3).
Proposition 4.2.
Let $L$ and $c_{1}$ be fixed positive constants. We consider a priori $a<0$ in (1.8) which determines the kink-soliton profile $\phi_{2,a}$ (therefore $\phi_{2,a}(L)>\pi$). Suppose that $\phi_{1}$ satisfies equation in (4.3) and $\phi_{1}(-L)=\phi_{1}(L)=\phi_{2,a}(L)$. Then, there is a family of solutions $\phi_{1}\equiv\phi_{1,k}$ for equation in (4.3) of subluminal type with a single-lobe profile on $[-L,L]$ given by,
$\phi_{1,k}(x)=2\pi-\arccos\Big{[}-1+2k^{2}\operatorname{sn}^{2}\Big{(}\frac{x}{c_{1}}+K(k);k\Big{)}\Big{]},\quad x\in[-L,L],$ (4.7)
with modulus $k$ satisfying $K(k)>\frac{L}{c_{1}}$. More precisely, for $\frac{L}{c_{1}}\leqq\frac{\pi}{2}$, we have $k\in(0,1)$, and for $\frac{L}{c_{1}}>\frac{\pi}{2}$, we have $k\in(k_{0},1)$, with $K(k_{0})=\frac{L}{c_{1}}$. The profile $\phi_{1,k}$ satisfies $\phi^{\prime\prime}_{1,k}(x)<0$ for all $x\in[-L,L]$.
Proof.
Let $a<0$. From the boundary condition $\phi_{1}(L)=\phi_{2,a}(L)>\pi$, we have that for all $x\in[-L,L]$, $2\pi\geqq\phi_{1}(0)\geqq\phi_{1}(x)\geqq\phi_{1}(L)>\pi$. Thus, from (4.6) we get the subluminal profile in (4.7).
Next, from condition $\phi_{1}(x)>\pi$ on $[-L,L]$, the property $\operatorname{sn}(2K(k);k)=0$ and (4.6), we need to have $\frac{x}{c_{1}}+K<2K$ for all $x\in[0,L]$ (see Figure 4 above). Therefore, a priori we get that modulus $k$ satisfies $K(k)>\frac{L}{c_{1}}$. Next, for $\frac{L}{c_{1}}\leqq\frac{\pi}{2}$ we obtain immediate a family of single-lobe solutions $k\to\phi_{1,k}$ for any $k\in(0,1)$. For $\frac{L}{c_{1}}>\frac{\pi}{2}$, we consider $k_{0}\in(0,1)$ such that $K(k_{0})=\frac{L}{c_{1}}$, then we have the family of single-lobe profiles $k\in(k_{0},1)\to\phi_{1,k}$ defined in (4.7). Lastly, from (4.3) we get immediately $\phi^{\prime\prime}_{1,k}(x)<0$ for all $x\in[-L,L]$. ∎
Remark 4.3.
We consider $L=\pi$, $c_{1}=1$, then $K(k)=\pi$ for $k\approx 0.9843\equiv k_{0}$, and so for $k>k_{0}$ we get the family of subluminal solutions $k\to\phi_{1,k}$ in (4.7) with a single-lobe profile on $[-L,L]$. Figure 4 shows the profile of $\phi_{1,k}$ for $k^{2}=0.98$.
Proposition 4.4.
Let $L,c_{1}$ be fixed positive constants. We consider a priori $a>0$ in (1.8) which determines the kink-soliton profile $\phi_{2,a}$ (with $\phi_{2,a}(L)<\pi$). Suppose that $\phi_{1}$ satisfies equation in (4.3) and $\phi_{1}(-L)=\phi_{1}(L)=\phi_{2,a}(L)$. Then, there is a family of solutions $\phi_{1}\equiv\phi_{1,k}$ for equation in (4.3) of subluminal type with a single-lobe profile on $[-L,L]$ given by
$\phi_{1,k}(x)=\left\{\begin{array}[]{ll}2\pi-\arccos\Big{[}-1+2k^{2}\operatorname{sn}^{2}\Big{(}\frac{x}{c_{1}}+K;k\Big{)}\Big{]},\quad\quad x\in[0,b],\\ \arccos\Big{[}-1+2k^{2}\operatorname{sn}^{2}\Big{(}\frac{x}{c_{1}}+K;k\Big{)}\Big{]},\;\qquad\quad\quad x\in[b,L],\end{array}\right.$ (4.8)
with $b=b(k)$ being unique in $(0,L)$ such that $\phi_{1,k}(b)=\pi$ (indeed, $b(k)=c_{1}K(k)$), and with modulus $k$ satisfying $K(k)<\frac{L}{c_{1}}$. Therefore, we have,
  1. (a)
    for $\frac{L}{c_{1}}\leqq\frac{\pi}{2}$ there are no single-lobe profiles for (4.3) with the profile in (4.8),
  2. (b)
    for $\frac{L}{c_{1}}>\frac{\pi}{2}$, $\phi_{1,k}$ in (4.8) is a single-lobe profile for (4.3) with $k\in(0,k_{0})$, and $k_{0}$ satisfying $K(k_{0})=\frac{L}{c_{1}}$; and,
  3. (c)
    $\phi^{\prime\prime}_{1,k}(b)=0$, $\phi^{\prime\prime}_{1,k}(x)<0$ for $x\in[-b,b]$ and $\phi^{\prime\prime}_{1,k}(x)>0$ for $x\in[-L,L]/[-b,b]$.
Proof.
From $a>0$, $\phi_{1}(0)>\pi$ and the boundary condition $\phi_{1}(L)=\phi_{2,a}(L)<\pi$, we have a unique $b\in(0,L)$, $b=b(k)$, such that $\phi_{1}(b)=\pi$, and so for $x\in[0,b]$, $\phi_{1}(x)\in[\pi,\phi_{1}(0)]$, and for $x\in[b,L]$, $\phi_{1}(x)\in[\phi_{1}(L),\pi]$. Thus, from (4.6) we get the single-lobe profile in (4.8). Next, from $\phi_{1,k}(b)=\pi$ and the strictly-decreasing property of $\phi_{1,k}$ on $[0,L]$ we need to have $b(k)=c_{1}K(k)$ (see Figure 7 above). Therefore, $K(k)=\frac{b}{c_{1}}<\frac{L}{c_{1}}$. Moreover, from (4.3) we obtain $\phi^{\prime\prime}_{1}(b)=0$, $\phi^{\prime\prime}_{1}(x)<0$ for $x\in[0,b)$ and $\phi^{\prime\prime}_{1}(x)>0$ for $x\in[b,L]$ (see Figure 9 below). ∎
Remark 4.5.
We consider $L=\pi$, $c_{1}=1$, then $K(k)=\pi$ for $k\approx 0.984432\equiv k_{0}$, and so for $k\in(0,k_{0})$ we get the family of subluminal librational solutions $k\to\phi_{1,k}$ in (4.8) with a single-lobe profile on $[-L,L]$ with $b=K(k)$. Figures 8 and 8 show the parts of $\phi_{1,k}$ in (4.8) when they are considered on the complete interval $[0,\pi]$ with $k=0.5$. Figure 9 shows the full profile of $\phi_{1,k}$ in (4.8) with $k=0.5$, $b=K(0.5)$ and it which is based on gluing the two components for $\phi_{1,k}$ in (4.8) from the profiles in Figures 8 and 8. We note that the “corners” (namely, the points where the derivative does not exist) in Figures 8 and 8 occur simultaneously at $x=b=K(0.5)$.
Figure 8: Graph of $\phi_{1,k}$ in (4.8) with $c_{1}=1$, $k=0.5$, $L=\pi$ and $x\in[0,\pi]$. Here $b\approx 1.85407\in(0,\pi)$. Panel (a) shows the first part with solid blue line for $x\in[0,b]$. Panel (b) shows the second part with solid blue line for $x\in[b,\pi]$ (color online).
Figure 9: Graph of the single-lobe solution (in solid blue line), composed by the extension of $\phi_{1,k}$ in (4.8) with $c_{1}=1$, $k=0.5$ to the whole interval $x\in[-\pi,\pi]$. The glueing of both solutions occurs at $x=b$ and $x=-b$ with $b\approx 1.85407$ (color online).
Proposition 4.6.
Let $L,c_{1},c_{2}$ be fixed positive constants. We consider $a=0$ in (1.8). Then, the degenerate single-lobe kink state $(\pi,\phi_{2,0})$ belongs to $D_{Z}$ if and only if $Z>0$ with $Z=\frac{2}{\pi c_{2}}$ (see Figure 10).
Proof.
From $a=0$ and $\phi_{1}(L)=\phi_{2,0}(L)=\pi$ we get that $\phi_{1}(x)\equiv\pi$ is the center solution for (4.3). Thus, $(\pi,\phi_{2,0})\in D_{Z}$ if and only if $Z>0$ with $Z=\frac{2}{\pi c_{2}}$. ∎
Figure 10: Degenerate single-lobe kink state $(\pi,\phi_{2,0})$ for the sine-Gordon model.
4.3 Existence of a single-lobe kink state
In this subsection we determine conditions for the profile $\Theta=(\phi_{1},\phi_{2})$ with $\phi_{2}$ in (1.8) and $\phi_{1}$ in (4.7) or (4.8) belongs to the $\delta$-interaction domain $D_{Z}$ in (1.7). Thus, from Propositions 4.2 and 4.4 we need to determine one expression for the shift-value $a=a(k)$ associated to the kink-profile $\phi_{2}=\phi_{2,a}$. From Proposition 4.1, we know that a priori the strength $Z$ satisfies $Z\in(-\infty,\frac{2}{\pi c_{2}})$.
Proposition 4.7.
Let $L,c_{1}$ fixed positive constants. For $a<0$ we consider the kink-soliton profile $\phi_{2,a}$ in (1.8) with $\phi_{2,a}(L)>\pi$ and the single-lobe profile $\phi_{1,k}$ in (4.7). We have a single-lobe kink state, namely $(\phi_{1,k},\phi_{2,a})\in D_{Z}$, in the following two cases:
  1. (1)
    Suppose that $\frac{L}{c_{1}}>\frac{\pi}{2}$. Then for $k_{0}\in(0,1)$ such that $K(k_{0})=\frac{L}{c_{1}}$, we require,
    1. (i)
      for $Z>0$ and $\frac{c_{1}}{2c_{2}}\geqq\tanh(\frac{L}{c_{1}})$ we need to have $Z\in\big{(}0,\frac{4}{\pi c_{1}}\big{[}\frac{c_{1}}{2c_{2}}-k_{0}\big{]}\big{)}$;
    2. (ii)
      for $Z>0$ and $k_{0}<\frac{c_{1}}{2c_{2}}<\tanh(\frac{L}{c_{1}})$ we need to have $Z\in\big{(}0,\frac{4}{\pi c_{1}}\big{[}\frac{c_{1}}{2c_{2}}-k_{0}\big{]}\big{)}$;
    3. (iii)
      for $Z<0$ and $\frac{c_{1}}{2c_{2}}<k_{0}$ we need to have $Z\in\big{(}\frac{4}{\pi c_{1}}\big{[}\frac{c_{1}}{2c_{2}}-k_{0}\big{]},0\big{)}$;
    4. (iv)
      for $Z<0$ and $\frac{c_{1}}{2c_{2}}=k_{0}$ there is $m_{0}<0$ such that we need to have $Z\in(m_{0},0)$;
    5. (v)
      for $Z<0$ and $k_{0}<\frac{c_{1}}{2c_{2}}<\tanh(\frac{L}{c_{1}})$ there is $m_{1}<0$ such that we need to have $Z\in(m_{1},0)$; and,
    6. (vi)
      for $Z=0$ there is a single-lobe kink state if and only if $\frac{c_{1}}{2c_{2}}\in(k_{0},\tanh(L/c_{1}))$.
  2. (2)
    Suppose that $\frac{L}{c_{1}}\leqq\frac{\pi}{2}$. Then we require,
    1. (i)
      $Z\in(0,\frac{2}{\pi c_{2}})$;
    2. (ii)
      for $Z<0$ and $\frac{c_{1}}{2c_{2}}<\tanh(\frac{L}{c_{1}})$ there is $m_{2}<0$ such that we need to have $Z\in(m_{2},0)$;
    3. (iii)
      $Z=0$ and $\frac{c_{1}}{2c_{2}}<\tanh(\frac{L}{c_{1}})$.
There is not single-lobe kink state, namely $(\phi_{1,k},\phi_{2,a})\notin D_{Z}$, in the following two cases:
  1. (3)
    Suppose that $\frac{L}{c_{1}}>\frac{\pi}{2}$. Then for $k_{0}\in(0,1)$ such that $K(k_{0})=\frac{L}{c_{1}}$ we have the following conditions:
    1. (i)
      for $Z>0$ and $\frac{c_{1}}{2c_{2}}\leqq k_{0}$ there is not a single-lobe kink state,
    2. (ii)
      for $Z<0$ and $\frac{c_{1}}{2c_{2}}\geqq\tanh(\frac{L}{c_{1}})$ there is not a single-lobe kink state,
  2. (4)
    Suppose $\frac{L}{c_{1}}\leqq\frac{\pi}{2}$. Then we have the following conditions:
    1. (i)
      for $Z<0$ and $\frac{c_{1}}{2c_{2}}\geqq\tanh(\frac{L}{c_{1}})$ there is not a single-lobe kink state,
    2. (ii)
      for $Z=0$ and $\frac{c_{1}}{2c_{2}}\geqq\tanh(\frac{L}{c_{1}})$ there is not a single-lobe kink state.
Remark 4.8.
The proof of Proposition 4.7 below shows formulae for $m_{0},m_{1},m_{2}$. Since $\tanh(K(k))>k$ for all $k\in(0,1)$, then in the case $(1)(i)$ above we necessarily have $\frac{c_{1}}{2c_{2}}-k_{0}>0$.
Proof of Proposition 4.7.
Let $a<0$. The idea in the following is to glue the subluminal-profile $\phi_{1,k}$ in (4.7) with $\phi=\phi_{2,a}$ at the vertex $\nu=L$ such that $\phi_{1,k}(L)=\phi_{2,a}(L)$ and with $k$ satisfying a priori $K(k)>\frac{L}{c_{1}}$.
Indeed, by formulae (1.8) and (4.7) we obtain the following first formula for the shift parameter $a=a(k)$,
$e^{-\frac{a}{c_{2}}}=\frac{\mathrm{dn}(L/c_{1};k)+k\mathrm{cn}(L/c_{1};k)}{k^{\prime}},$ (4.9)
where $\mathrm{dn}(\cdot;k)$ and $\mathrm{cn}(\cdot;k)$ represent the Jacobian elliptic functions of dnoidal and cnoidal type, respectively (see Figure 7). Here, $k^{\prime}$ is defined by $k^{\prime}=\sqrt{1-k^{2}}$. In the following, we show formula in (4.9). Indeed, for $\beta\equiv\arctan(e^{-\frac{a}{c_{2}}})$ we get from (4.6) that $\cos(4\beta)=-1+2k^{2}sn^{2}\Big{(}\frac{L}{c_{1}}+K;k\Big{)}$. Then, from $\cos(4\beta)=2\cos^{2}(2\beta)-1$ follows $k^{2}\operatorname{sn}^{2}(\frac{L}{c_{1}}+K;k)=\cos^{2}(2\beta)$ Thus, since $2\beta\in[\frac{\pi}{2},\pi]$ and $\frac{L}{c_{1}}+K<2K$ we get $\cos(2\beta)=-k\operatorname{sn}(\frac{L}{c_{1}}+K;k)$ (see, for instance, Figure 4). Moreover, from $\cos^{2}(2\beta)+\sin^{2}(2\beta)=1$ and $1-k^{2}\operatorname{sn}^{2}=\mathrm{dn}^{2}$ we also obtain $\sin(2\beta)=\mathrm{dn}(\frac{L}{c_{1}}+K;k)$. Therefore,
$e^{-\frac{a}{c_{2}}}=\tan(\beta)=\frac{1-\cos(2\beta)}{\sin(2\beta)}=\frac{1+k\operatorname{sn}(\frac{L}{c_{1}}+K;k)}{\mathrm{dn}(\frac{L}{c_{1}}+K;k)}=\frac{\mathrm{dn}(L/c_{1};k)+k\mathrm{cn}(L/c_{1};k)}{k^{\prime}},$ (4.10)
where we are using the relations $\operatorname{sn}(u+K)=\mathrm{cn}(u)/\mathrm{dn}(u)$ and $\mathrm{dn}(u+K)=k^{\prime}/\mathrm{dn}(u)$ (cf. Byrd and Friedman [22]).
Now, from (4.9) we obtain the following more manageable relationship between $a=a(k)$ and $k$,
$\operatorname{sech}\Big{(}\frac{a(k)}{c_{2}}\Big{)}=\frac{k^{\prime}}{\mathrm{dn}(L/c_{1};k)},$ (4.11)
and so we get the property that the shift-value mapping $k\to a(k)$ will be smooth. Formula (4.11) shows that always we can glue a librational profile with a kink-profile because $k^{\prime}\leqq\mathrm{dn}(x;k)$ for all $x$ and $k$.
In the sequel we determine the correct values of $k$ and $Z$ such that $(\phi_{1,k},\phi_{2,a(k)})\in D_{Z}$, namely, such that we have the relation $2\phi^{\prime}_{1}(L)=\phi^{\prime}_{2}(L)+Z\phi_{2}(L)$. We divide this study into several cases.
$\bullet$ Case (1). Let $\frac{L}{c_{1}}>\frac{\pi}{2}$. Consider the unique $k_{0}\in(0,1)$ such that $K(k_{0})=\frac{L}{c_{1}}$. Then for $k>k_{0}$ it follows that $K(k)>\frac{L}{c_{1}}$. Now, from properties of the Jacobi elliptic functions, we obtain from (4.11) the following limits:
$\lim_{k\to 1}\operatorname{sech}\Big{(}\frac{a(k)}{c_{2}}\Big{)}$ $=\lim_{k\to 1}\frac{k^{\prime}}{\mathrm{dn}(L/c_{1};k)}=\frac{0}{\operatorname{sech}(L/c_{1})}=0,$ (4.12)
$\lim_{k\to k_{0}}\operatorname{sech}\Big{(}\frac{a(k)}{c_{2}}\Big{)}$ $=\frac{k^{\prime}_{0}}{\mathrm{dn}(K(k_{0});k_{0})}=\frac{k^{\prime}_{0}}{k^{\prime}_{0}}=1.$
Therefore, $\lim_{k\to 1}a(k)=-\infty$ and $\lim_{k\to k_{0}}a(k)=0$ ($a(k_{0})=0$). Next, we will see that $k\in(k_{0},1)\to a(k)$ is a strictly decreasing mapping. Indeed, from the formula $\operatorname{sech}^{-1}(x)=\ln(\frac{1+\sqrt{1-x^{2}}}{x})$, $0<x\leqq 1$, and $\mathrm{dn}^{2}-k^{\prime 2}=k^{2}\mathrm{cn}^{2}$ we get the following relation
$\frac{a(k)}{c_{2}}=-\ln\Big{(}\frac{\mathrm{dn}(L/{c_{1}};k)+k\mathrm{cn}(L/{c_{1}};k)}{k^{\prime}}\Big{)}\equiv j(k).$ (4.13)
Then, numerical simulations always show that $j(k)$ in (4.13) is strictly decreasing for $k\in(k_{0},1)$ and so $a^{\prime}(k)<0$, for $k\in(k_{0},1)$ (see Figure 11 for the graph of $j=j(k)$, $k\in(k_{0},1)$ with $c_{1}=1$, $L=\pi$ and $k_{0}\approx 0.984432$). We note that by using formula (710.53) in [22] and deriving with regard to $k$ both sides of equality in (4.11) we can also see that $a(k)$ is strictly decreasing after long calculations.
Figure 11: Graph of the function $j=j(k)$ in (4.13) for $k\in(k_{0},1)$, with $c_{1}=1$, $L=\pi$ and where $k_{0}\approx 0.984432$.
Next, from (4.4), $\operatorname{sn}^{2}+\mathrm{cn}^{2}=1$, $\frac{L}{c_{1}}+K\in[K,2K]$ and $\mathrm{cn}(u+K)=-k^{\prime}\operatorname{sn}(u)/\mathrm{dn}(u)$ we get for $\phi_{1}=\phi_{1,k}$,
$\phi^{\prime}_{1}(L)=\frac{2k}{c_{1}}\mathrm{cn}(L/c_{1}+K;k)=-\frac{2kk^{\prime}\operatorname{sn}(L/c_{1};k)}{c_{1}\mathrm{dn}(L/c_{1};k)}.$ (4.14)
From (1.8) and (4.11) we get for $\phi_{2}=\phi_{2,a}$,
$\phi^{\prime}_{2}(L)=-\frac{2}{c_{2}}\operatorname{sech}\Big{(}\frac{a}{c_{2}}\Big{)}=-\frac{2}{c_{2}}\frac{k^{\prime}}{\mathrm{dn}(L/c_{1};k)}.$ (4.15)
Thus, we obtain from $2\phi^{\prime}_{1}(L)=\phi^{\prime}_{2}(L)+Z\phi_{2}(L)$ the following relation
$\frac{k^{\prime}}{c_{1}\mathrm{dn}(L/c_{1};k)}\Big{[}\frac{c_{1}}{2c_{2}}-k\operatorname{sn}(L/c_{1};k)\Big{]}=Z\beta,$ (4.16)
with $\beta\equiv\arctan(e^{-\frac{a}{c_{2}}})$. Therefore for $a=a(k)<0$ and $k\in(k_{0},1)$, $Z$ and $k$ are related by the equation
$Z=\frac{\operatorname{sech}(a/c_{2})}{c_{1}\arctan(e^{-a/c_{2}})}\Big{[}\frac{c_{1}}{2c_{2}}-k\operatorname{sn}(L/c_{1};k)\Big{]}\equiv H(k).$ (4.17)
Thus, the existence of single-lobe kink profile for the sine-Gordon model on a tadpole is reduced to solve the equation
$H(k)=Z,$ (4.18)
for $Z\in(-\infty,\frac{2}{\pi c_{2}})$ fixed, and some $k\in(k_{0},1)$ with $a=a(k)$ solving (4.11). Next, we consider the following cases:
  1. (I)
    Let $Z>0$. Then, a priori, $k$ needs to satisfy $k\operatorname{sn}(L/c_{1};k)<\frac{c_{1}}{2c_{2}}$ for solving (4.18). Thus, from basic properties of the Jacobi elliptic functions we get the following limits
    $\lim_{k\to 1}k\operatorname{sn}(L/c_{1};k)$ $=\tanh(L/c_{1}),$ (4.19)
    $\lim_{k\to k_{0}}k\operatorname{sn}(L/c_{1};k)$ $=k_{0}\operatorname{sn}(K(k_{0});k_{0})=k_{0}.$
    Moreover, the mapping
    $k\in(k_{0},1)\to g(k)\equiv k\operatorname{sn}(L/c_{1};k),$ (4.20)
    is strictly increasing (we note that for $k<k_{0}$ it is possible to have $g(k)\leqq 0$ and with $g$ having some oscillations; see Figures 12 and 13), and so we get
    $k_{0}=k_{0}\operatorname{sn}(K(k_{0});k_{0})<k\operatorname{sn}(L/c_{1};k)<\tanh(L/c_{1}),\;\;\text{for}\;\;k\in(k_{0},1).$
    Figure 12: Graph of the function $g(k)$ defined on (4.20) for $k\in(0,1)$, with $c_{1}=1$, $L=2.5\pi$.
    Figure 13: Graph of $g(k)$ for $k\in(0.99,1)$, with $c_{1}=1$, $L=2.5\pi$ and $k^{2}_{0}\approx 0.999998$.
    Let us consider the following subcases:
    1. (i)
      Suppose $\frac{c_{1}}{2c_{2}}\geqq\tanh(L/c_{1})$. Then $H(k)>0$ for $k\in(k_{0},1)$. Now, from (4.12) and properties of the Jacobi elliptic functions (cf. [22]), we obtain
      $\lim_{k\to 1}H(k)$ $=\frac{0}{\arctan(\infty)}\Big{[}\frac{c_{1}}{2c_{2}}-\tanh(L/c_{1})\Big{]}=0,$ (4.21)
      $\lim_{k\to k_{0}}H(k)$ $=\frac{4}{\pi c_{1}}\Big{[}\frac{c_{1}}{2c_{2}}-k_{0}\operatorname{sn}(K(k_{0});k_{0})\Big{]}=\frac{4}{\pi c_{1}}\Big{[}\frac{c_{1}}{2c_{2}}-k_{0}\Big{]}>0.$
      Thus, by analysis above ($a^{\prime}(k)<0$ and $g^{\prime}(k)>0$) we can see that $H^{\prime}(k)<0$ for $k\in(k_{0},1)$, and therefore for $Z\in\big{(}0,\frac{4}{\pi c_{1}}\big{[}\frac{c_{1}}{2c_{2}}-k_{0}\big{]}\big{)}$ there is a unique $k_{Z}\in(k_{0},1)$ with $H(k_{Z})=Z$. This shows $(1)(i)$.
    2. (ii)
      Suppose $k_{0}<\frac{c_{1}}{2c_{2}}<\tanh(L/c_{1})$. Then, there is a unique $\gamma\in(k_{0},1)$ with $g(\gamma)=\frac{c_{1}}{2c_{2}}$ and such that for $k\in(k_{0},\gamma)$ follows $H(k)>0$. From $\lim_{k\to\beta}H(k)=0$ and (4.21), we have for $Z\in\big{(}0,\frac{4}{\pi c_{1}}\big{[}\frac{c_{1}}{2c_{2}}-k_{0}\big{]}\big{)}$ that there is a unique $k_{Z}\in(k_{0},\beta)$ with $H(k_{Z})=Z$. This shows $(1)(ii)$.
  2. (II)
    Now let $Z<0$. From (4.17) follows that $k$ a priori needs to satisfy $\frac{c_{1}}{2c_{2}}<k\operatorname{sn}(L/c_{1};k)<\tanh(L/c_{1})$. Let us examine the following subcases:
    1. (iii)
      Suppose $\frac{c_{1}}{2c_{2}}<k_{0}$. Then $g(k)>k_{0}$ for $k\in(k_{0},1)$, and so $H(k)<0$ and $H^{\prime}(k)>0$ for any $k\in(k_{0},1)$. Therefore, for any $Z\in\big{(}\frac{4}{\pi c_{1}}\big{[}\frac{c_{1}}{2c_{2}}-k_{0}\big{]},0\big{)}$ there is a unique $k_{Z}\in(k_{0},1)$ with $H(k_{Z})=Z$. This shows $(1)(iii)$.
    2. (iv)
      Let $\frac{c_{1}}{2c_{2}}=k_{0}$. Then $\lim_{k\to k_{0}}H(k)=0$. Let $m_{0}=\min_{k\in(k_{0},1)}H(k)$, then for $Z\in[m_{0},0)$ there is at least $k_{Z}\in(k_{0},1)$ such that $H(k_{Z})=Z$. This proves $(1)(iv)$.
    3. (v)
      Next, for $k_{0}<\frac{c_{1}}{2c_{2}}<\tanh(L/c_{1})$ we consider $\beta\in(k_{0},1)$ with $g(\beta)=\frac{c_{1}}{2c_{2}}$. Then, for every $k\in(\beta,1)$ follows $g(k)>\frac{c_{1}}{2c_{2}}$ and so $H(k)<0$ for $k\in(\beta,1)$. Since $\lim_{k\to\beta}H(k)=0$, we obtain for $Z\in(m_{\beta},0)$, $m_{\beta}=\min_{k\in(\beta,1)}H(k)$, that there is at least $k_{Z}\in(\beta,1)$ such that $H(k_{Z})=Z$. This shows $(1)(v)$
  3. (III)
    Now suppose that $Z=0$. From (4.17) we need to have $H(k)=0$ and so $\frac{c_{1}}{2c_{2}}=k\operatorname{sn}(L/c_{1};k)$ is true if and only if $\frac{c_{1}}{2c_{2}}\in(k_{0},\tanh(L/c_{1}))$ and in this case there is a unique $k\in(k_{0},1)$ such that $H(k)=0$. This shows $(1)(vi)$.
$\bullet$ Case (2). Now consider the case $\frac{L}{c_{1}}\leqq\frac{\pi}{2}$. Then $K(k)>\frac{L}{c_{1}}$ for all $k\in(0,1)$. Thus, since $\mathrm{dn}(\cdot;0)=1$ we get from (4.11) and (4.12) again $\lim_{k\to 1}a(k)=-\infty$ and $\lim_{k\to 0}a(k)=0$. Moreover, the mapping $k\in(0,1)\to g(k)=k\operatorname{sn}(L/c_{1};k)$ is strictly increasing (see Figure 13) with $k\operatorname{sn}(L/c_{1};k)>0$, $\lim_{k\to 1}k\operatorname{sn}(L/c_{1};k)=\tanh(\frac{L}{c_{1}})$ and $\lim_{k\to 0}k\operatorname{sn}(L/c_{1};k)=0$.
Figure 14: Graph of $G(k,x)=k\operatorname{sn}(x;k)$ for $k\in(0,1)$ and $x\in(0,\frac{\pi}{2})$.
Once again, we split the analysis into subcases:
  1. (I)
    Let $Z>0$. Then we have:
    1. (i)
      For $\frac{c_{1}}{2c_{2}}\geqq\tanh(\frac{L}{c_{1}})>k\operatorname{sn}(L/c_{1};k)$ it follows that $H(k)>0$ for $k\in(0,1)$ and $\lim_{k\to 0}H(k)=\frac{2}{\pi c_{2}}$, $\lim_{k\to 1}H(k)=0$, and $H^{\prime}(k)<0$ for all $k$. Then, for every $Z\in(0,\frac{2}{\pi c_{2}})$ there is a unique $k_{Z}\in(0,1)$ with $H(k_{Z})=Z$.
    2. (ii)
      For $\frac{c_{1}}{2c_{2}}<\tanh(\frac{L}{c_{1}})$ consider $\beta\in(0,1)$ with $g(\beta)=\frac{c_{1}}{2c_{2}}$. Then, for all $k\in(0,\beta)$ we obtain $H(k)>0$. From $\lim_{k\to 0}H(k)=\frac{2}{\pi c_{2}}$ and $\lim_{k\to\beta}H(k)=0$, we obtain for every $Z\in(0,\frac{2}{\pi c_{2}})$ that there is a unique $k_{Z}\in(0,\beta)$ with $H(k_{Z})=Z$.
    This proves $(2)(i)$.
  2. (II)
    Let $Z<0$. For $\frac{c_{1}}{2c_{2}}<\tanh(\frac{L}{c_{1}})$ we consider $\beta\in(0,1)$ with $g(\beta)=\frac{c_{1}}{2c_{2}}$, then $H(k)<0$ for all $k\in(\beta,1)$. From $\lim_{k\to 1}H(k)=0$ and $\lim_{k\to\beta}H(k)=0$, for $Z\in(p_{\beta},0)$, $p_{\beta}=\min_{k\in(\beta,1)}H(k)$, there is at least $k_{Z}\in(\beta,1)$ such that $H(k_{Z})=Z$. This proves $(2)(ii)$.
  3. (III)
    Let $Z=0$. For $\frac{c_{1}}{2c_{2}}<\tanh(\frac{L}{c_{1}})$, we consider the unique $k\in(0,1)$ with $g(k)=\frac{c_{1}}{2c_{2}}$. Then $H(k)=0$. This shows $(2)(iii)$.
$\bullet$ Case (3). Suppose $\frac{L}{c_{1}}>\frac{\pi}{2}$ and we consider $k_{0}\in(0,1)$ such that $K(k_{0})=\frac{L}{c_{1}}$.
  1. (i)
    If $Z>0$ and $\frac{c_{1}}{2c_{2}}\leqq k_{0}$ then $\frac{c_{1}}{2c_{2}}-g(k)<0$ and so we do not haven a solution for $H(k)=Z$. Therefore there is not a single-lobe kink state. This proves $(3)(i)$.
  2. (ii)
    If $Z<0$ and $\frac{c_{1}}{2c_{2}}\geqq\tanh(\frac{L}{c_{1}})$ then $\frac{c_{1}}{2c_{2}}-g(k)>0$ and obviously there is not a single-lobe kink state. This shows $(3)(ii)$.
$\bullet$ Case (4). Finally, suppose that $\frac{L}{c_{1}}\leqq\frac{\pi}{2}$. Then we have the following conditions:
  1. (i)
    For $Z<0$ and $\frac{c_{1}}{2c_{2}}\geqq\tanh(\frac{L}{c_{1}})$ we have $\frac{c_{1}}{2c_{2}}-g(k)>0$ and, clearly, there is not a single-lobe kink state. This proves $(4)(i)$
  2. ii)
    For $Z=0$ and $\frac{c_{1}}{2c_{2}}\geqq\tanh(\frac{L}{c_{1}})$ there holds $\frac{c_{1}}{2c_{2}}-g(k)>0$ and, clearly, there is not a single-lobe kink state. This shows $(4)(ii)$.
The Proposition is now proved. ∎
In the sequel we formally prove that, for the Neumann-Kirchhoff condition in (1.7), namely, $Z=0$, we get a single-lobe kink state for the sine-Gordon model with a profile on $[-L,L]$ given by formula in (4.8). For $Z\neq 0$ we found technical difficulties in determining the existence of these profiles (see Remark 4.10).
Proposition 4.9.
Let $L,c_{1}$ be fixed positive constants with $\frac{L}{c_{1}}>\frac{\pi}{2}$. For the subluminal profile $\phi_{1,k}$ in (4.8) with $k\in(0,k_{0})$ and $K(k_{0})=\frac{L}{c_{1}}$, there is a shift-value $a=a(k)>0$ associated to the kink-profile $\phi_{2,a}$ in (1.8) such that $(\phi_{1,k},\phi_{2,a})\in D_{Z}$, with $Z=0$, if and only if $c_{2}$ satisfies
$\frac{c_{1}}{2k_{0}}<c_{2}.$
Moreover, $a(k)>0$ is given univocally by relation in (4.11). For $Z<0$, there is not a single-lobe kink state with a first-component profile $\phi_{1,k}$ in (4.8) with $c_{2}$ satisfying $\frac{c_{1}}{2c_{2}}\geqq k_{0}$.
Proof.
We show initially that for $x\in(b,L]$, $2K<\frac{x}{c_{1}}+K<3K$ for every admissible $k$. Indeed, suppose $\beta\in(b,L]$ such that $2K\geqq\frac{\beta}{c_{1}}+K$ then since $b=c_{1}K$ follows immediately $2K>\frac{b}{c_{1}}+K=2K$. Next, for all $k\in(0,k_{0})$ we have $\frac{L}{c_{1}}+K(k)<3K(k)$, in fact, suppose there is $p\in(0,k_{0})$ such that $\frac{L}{c_{1}}+K(p)>3K(p)$ then there is $\beta\in(0,p)$ satisfying $\frac{L}{c_{1}}+K(\beta)=3K(\beta)$. Thus, from relations (4.4) and (4.8) we get
$-\frac{c_{1}^{2}}{2}[\phi_{1,\beta}^{\prime}(L)]^{2}=\cos(\phi_{1,\beta}(L))-\cos(\phi_{1,\beta}(0))=-2\beta^{2}\mathrm{cn}^{2}\Big{(}\frac{L}{c_{1}}+K(\beta);\beta\Big{)}=0.$ (4.22)
Therefore, $\phi_{1,\beta}^{\prime}(L)=0$, which is clearly not possible. We note from the profile of the elliptic function $\operatorname{sn}(\cdot)$ that $-1\leqq\operatorname{sn}(\frac{x}{c_{1}}+K)\leqq 0$ for $x\in(b,L]$ and for $x=L$, $K<\frac{L}{c_{1}}<2K$ and hence $\operatorname{sn}(\frac{L}{c_{1}})>0$ and $\mathrm{cn}(\frac{L}{c_{1}})<0$.
Next, to show the continuity condition $\phi_{1,k}(L)=\phi_{2,a}(L)$ for $a=a(k)>0$ we shall follow the same strategy as in the proof of Proposition 4.7. Indeed, we get the first same expression for the shift $a$ as in (4.9) by considering again $\beta\equiv\arctan(e^{-\frac{a}{c_{2}}})$ but with $\beta\in(0,\frac{\pi}{4}]$. Thus, from relation $k^{2}\operatorname{sn}^{2}(\frac{L}{c_{1}}+K;k)=\cos^{2}(2\beta)$ we get $\cos(2\beta)=-k\operatorname{sn}(\frac{L}{c_{1}}+K;k)$ and $\sin(2\beta)=\mathrm{dn}(\frac{L}{c_{1}}+K;k)$. Thus, we obtain (4.9) with $a>0$ and a similar relation as in (4.11). Next, we determine the right $k\in(0,k_{0})$ such that $2\phi_{1,k}^{\prime}(L)=\phi^{\prime}_{2,a(k)}(L)$. Initially we get
$\lim_{k\to 0}\operatorname{sech}\Big{(}\frac{a(k)}{c_{2}}\Big{)}=1,\quad\lim_{k\to k_{0}}\operatorname{sech}\Big{(}\frac{a(k)}{c_{2}}\Big{)}=1.$ (4.23)
Therefore, $\lim_{k\to 0}a(k)=0$ and $\lim_{k\to k_{0}}a(k)=0$ ($a(k_{0})=0$). We note that, compared to the negative case of $a$, $a=a(k)$ is no longer a strictly increasing function of $k$. In Figure 15 we show a generic profile of $a(k)$ given by formula in (4.13) and we notice that there is always a unique value $k_{\mathrm{crit}}\in(0,k_{0})$ such that $a^{\prime}(k_{\mathrm{crit}})=0$.
Figure 15: Graph of the function $a=a(k)$ for $k\in(0,k_{0})$, with $c_{1}=1$, $L=\pi$ and $k_{0}\approx 0.984433$, $K(k_{0})=\pi$.
Now, as relations for $\phi_{1,k}^{\prime}(L)$ and $\phi_{2,a}^{\prime}(L)$ in (4.14) and (4.15), respectively, holds true, we obtain $2\phi_{1,k}^{\prime}(L)=\phi^{\prime}_{2,a(k)}(L)$ if and only if $k$ satisfies the following relation
$k\operatorname{sn}\Big{(}\frac{L}{c_{1}};k\Big{)}=\frac{c_{1}}{2c_{2}}.$ (4.24)
We now study the mapping $k\in(0,k_{0})\to F(k)=k\operatorname{sn}\Big{(}K(k_{0});k\Big{)}$:
$F(0)=0,\quad\lim_{k\to k_{0}}F(k)=k_{0}\operatorname{sn}\Big{(}K(k_{0});k_{0}\Big{)}=k_{0},\quad F(k)\leqq k_{0}.$ (4.25)
Therefore, for $c_{2}$ satisfying $\frac{c_{1}}{2k_{0}}\geqq c_{2}$ there is not single-lobe kink state for the sine-Gordon model on a tadpole graph. Now, we analyze the case for $c_{2}$ satisfying $\frac{c_{1}}{2k_{0}}<c_{2}$. We note that in general the mapping $k\to F(k)$ has a profile with several oscillations (for instance, for large $L$) with finite simple zeros $\{0,r_{1},...,r_{n}\}\subset[0,k_{0})$ and for $k\in(r_{n},k_{0})$ this mapping is strictly increasing, so from (4.25) there is a unique $k_{c_{2}}\in(r_{n},k_{0})$ such that $F(k_{c_{2}})=\frac{c_{1}}{2c_{2}}$. Therefore, we obtain the existence of single-lobe kink profile $(\phi_{1,k_{c_{2}}},\phi_{2,a(k_{c_{2}})})\in D_{0}$ for the sine-Gordon model.
Let $Z<0$, then from equation $H(k)=Z$ in (4.18) with $H$ defined in (4.17) for $k\in(0,k_{0})$, we need to have a priori that $\frac{c_{1}}{2c_{2}}<k\operatorname{sn}(L/c_{1};k)$, for some $k\in(0,k_{0})$. Therefore, for $\frac{c_{1}}{2c_{2}}\geqq k_{0}$ there is not a single-lobe kink state. This finishes the proof. ∎
Remark 4.10.
A few observations are in order.
  1. (a)
    Let us illustrate the case $L=3\pi$, $c_{1}=1$, with $k_{0}$ satisfying $K(k_{0})=3\pi$ ($k_{0}\approx 1$). Figure 16 shows the profile of $F(k)=k\operatorname{sn}\big{(}K(k_{0});k\big{)}$ for $k\in(0,k_{0})$. Now, $F(k)=0$ if and only if $r_{1}\approx 0.84$ and $r_{2}\approx 0.9987$ (see Figures 17 and 17). By chosing $c_{2}=2$, we have that $F(k)=\frac{1}{4}$ if and only if $k_{1}\approx 0.32$, $k_{2}\approx 0.81$ and $k_{3}\approx 0.999$. But the right $k_{i}$ that produces a single-lobe kink profile ($\phi_{1,k}^{\prime}(x)<0$ for all $x\in[0,3\pi]$) will be exactly $k_{3}\approx 0.999$. In Figures 18, 18 and 18 we show the profile of $\phi_{1,k_{i}}(x)$ in (4.8) with $x\in[K(k_{i}),3\pi]$.
    Figure 16: $F(k)=k\operatorname{sn}(K(k_{0});k)$ for $k\in(0,k_{0})$, $K(k_{0})=3\pi$.
    Figure 17: $F(k)=k\operatorname{sn}(K(k_{0});k)$ for $k\in(0.8,k_{0})$ (panel (a)) and for $k\in(0.99,k_{0})$ (panel (b)); here $K(k_{0})=3\pi$. (color online).
    Figure 18: Graph of $\phi_{1,k}$ (in solid blue line) for $k=k_{1}$ (panel (a)), $k=k_{2}$ (panel (b)) and $k=k_{3}$ (panel (c)) in the domains $x\in[K(k_{i}),3\pi]$, $i=1,2,3$ (color online).
  2. (b)
    For the case $Z<0$ and $c_{2}$ such that $k_{0}>\frac{c_{1}}{2c_{2}}$, some technical problems arise in studying equation $H(k)=Z$ in (4.18) to obtain a general picture of existence of single-lobe kink profile as was that obtained in Proposition 4.7. The problem here is that the mapping $k\to k\operatorname{sn}(L/c_{1};k)$, $k\in(0,k_{0})$, may have a profile with several oscillations that do not allow us to clearly study the mapping $H$ in (4.17). A similar situation is obtained for $Z>0$.
5 Spectral structure associated to positive single-lobe kink states
Let us consider $\Theta=(\phi_{1},\phi_{2})\in D_{Z}$, an arbitrary positive single-lobe kink state for (1.6), and the family of self-adjoint operators $(\mathcal{L}_{Z},D_{Z})$,
$\mathcal{L}_{Z}=\Big{(}\Big{(}-c_{j}^{2}\frac{d^{2}}{dx^{2}}+\cos(\phi_{j})\Big{)}\delta_{j,k}\Big{)},\quad 1\leqq j,k\leqq 2,\;\;c_{j}>0.$ (5.1)
In this section we characterize the structure of the Morse and nullity indices of the diagonal Schrödinger operators $\mathcal{L}_{Z}$.
We start with the following “moving framework”. For $(f,g)\in D_{Z}$ consider $h(x)=g(x+L)$ for $x>0$. Then $h(0)=g(L)$ and $h^{\prime}(0)=g^{\prime}(L)$. Therefore, the eigenvalue problem $\mathcal{L}_{Z}(f,g)^{\top}=\lambda(f,g)^{\top}$ is equivalent to the following system
$\left\{\begin{array}[]{ll}\mathcal{L}_{1}f(x)=\lambda f(x),\quad\quad x\in(-L,L),\\ \mathcal{L}_{a}h(x)=\lambda h(x),\quad\quad x\in(0,\infty),\\ (f,h)\in D_{Z,0},\end{array}\right.$ (5.2)
where
$\mathcal{L}_{1}=-c_{1}^{2}\partial_{x}^{2}+\cos(\phi_{1}),\;\;\;\;\mathcal{L}_{a}\equiv-c_{2}^{2}\partial_{x}^{2}+\cos(\psi_{a})$ (5.3)
with
$\psi_{a}(x)=\phi_{2}(x+L)=4\arctan\Big{(}e^{-\frac{1}{c_{2}}(x+a)}\Big{)},\qquad x>0,$ (5.4)
representing the kink-soliton profile for any shift-value $a\in\mathbb{R}$. Here,
$D_{Z,0}=\big{\{}(f,h)\in X^{2}(-L,L):f(L)=f(-L)=h(0),\;\text{and},\;f^{\prime}(L)-f^{\prime}(-L)=h^{\prime}(0)+Zh(0)\big{\}}$ (5.5)
with $X^{n}(-L,L)\equiv H^{n}(-L,L)\oplus H^{n}(0,\infty)$, $n\in\mathbb{N}$. Naturally, $(\phi_{1},\psi_{a})\in D_{Z,0}$. For convenience of notation, we will use $\mathcal{L}\equiv(\mathcal{L}_{1},\mathcal{L}_{a})$.
In what follows, we establish our main strategy in order to study the eigenvalue problem in (5.2). More precisely, we will reduce the eigenvalue problem for $(\mathcal{L},D_{Z,0})$ into two classes of eigenvalue problems, one for $\mathcal{L}_{1}$ with periodic boundary conditions on $[-L,L]$, and one for the operator $\mathcal{L}_{a}$ with $\delta$-type boundary conditions on the half-line $(0,\infty)$. Thus, we employ the so-called splitting eigenvalue method on a tadpole graph (see Angulo [6, 7]).
Lemma 5.1 (splitting eigenvalue method).
Suppose that $(f,g)\in D_{Z,0}$ with $g(0)\neq 0$, satisfies the relation $\mathcal{L}(f,g)^{\top}=\lambda(f,g)^{\top}$ for $\lambda\leqq 0$. Then, we obtain the following two eigenvalues problems
$(PBP)\left\{\begin{array}[]{ll}\mathcal{L}_{1}f(x)=\lambda f(x),\quad x\in(-L,L),\\ f(-L)=f(L),\\ f^{\prime}(-L)=f^{\prime}(L),\end{array}\right.\quad(\delta BP)\left\{\begin{array}[]{ll}\mathcal{L}_{a}g(x)=\lambda g(x),\quad\quad x\in(0,\infty),\\ g^{\prime}(0+)=-Zg(0+).\end{array}\right.$ (5.6)
Proof.
Following the analysis by Angulo [6, 7], for $(f,g)\in D_{Z,0}$ and $g(0)\neq 0$, we have
$f(-L)=f(L),\;\;f^{\prime}(L)-f^{\prime}(-L)=\Big{[}\frac{g^{\prime}(0+)}{g(0+)}+Z\Big{]}f(L)\equiv\theta f(L).$
Now, we consider the following eigenvalue problem for $\mathcal{L}_{1}$ with real coupled self-adjoint boundary conditions determined by $\alpha\in\mathbb{R}$,
$(RC_{\alpha}):\;\;\left\{\begin{array}[]{ll}\mathcal{L}_{1}y(x)=\eta y(x),\quad\quad x\in(-L,L),\\ y(L)=y(-L),\\ y^{\prime}(L)-y^{\prime}(-L)=\alpha y(L).\end{array}\right.$ (5.7)
Thus, $f$ in (5.6) satisfies the real coupled problem $(RC_{\alpha})$ in (5.7) wth $\alpha=\theta$ and $\eta=\lambda$. We shall verify that $\theta=0$, which proves the Lemma. Indeed, we consider $K_{\theta}=[k_{ij}]$ the $2\times 2$-matrix associated to (5.7) given by $k_{11}=1$, $k_{12}=0$, $k_{21}=\theta$ and $k_{22}=1$ ($\text{det}(K_{\theta})=1$), and by $\eta_{n}=\eta_{n}(K_{\theta})$, $n\in\mathbb{N}_{0}$, the eigenvalues for the $(RC_{\theta})$-problem in (5.7). We also consider $\mu_{n}=\mu_{n}(K_{\theta})$ and $\nu_{n}=\nu_{n}(K_{\theta})$, $n\in\mathbb{N}_{0}$, the eigenvalues problems in (5.7) induced by $K_{\theta}$ with the following boundary conditions
[…]
respectively. We recall that if $y_{n}$ is an eigenfunction of $\mu_{n}$, then $y_{n}$ is unique up to constant multiples and has exactly $n$ zeros in $(-L,L)$, $n\in\mathbb{N}_{0}$ (a similar result is obtained for $u_{n}$ an eigenfunction of $\nu_{n}$, $n\in\mathbb{N}_{0}$). Next, by Theorem 4.8.1 in Zettl [52] we have that $\nu_{0}$ and $\eta_{0}$ are simple and, in particular, we have the following partial distribution of eigenvalues,
$\nu_{0}\leqq\eta_{0}<\{\mu_{0},\nu_{1}\}<\eta_{1}\leqq\{\mu_{1},\nu_{2}\}\leqq\eta_{2},$ (5.9)
where the notation $\{\mu_{n},\nu_{m}\}$ is used to indicate either $\mu_{n}$ or $\nu_{m}$ (but no comparison is made between $\mu_{n}$ and $\nu_{m}$). Next, we observe that $u^{\prime}_{0}(L)=0$ and so $\theta=0$ (because by (5.9) we have $\nu_{0}<\mu_{0}$ and so $u_{0}(\pm L)\neq 0$). Indeed, since $\nu_{0}$ is simple and the profile-solution $\phi_{1}$ is even, we need to have that $u_{0}$ is even or odd, but as $u_{0}$ has not zeros in $[-L,L]$ we get that $u_{0}$ is actually even. Thus, $u^{\prime}_{0}$ is odd and $u^{\prime}_{0}(\pm L)=0$. This finishes the proof. ∎
5.1 Morse index for $\mathcal{L}_{Z}$
The purpose of this subsection is to characterize the number of negative eigenvalues for $\mathcal{L}\equiv(\mathcal{L}_{1},\mathcal{L}_{a})$ with $\mathcal{L}_{1},\mathcal{L}_{a}$ defined in (5.3) and determined by a priori $(\phi_{1},\psi_{a})$ positive single-lobe kink state, $\psi_{a}$ in (5.4) and $a\in\mathbb{R}$. By the splitting eigenvalue result in Lemma 5.1, it arises naturally to study the Morse index for the family of self-adjoint operators,
$\mathcal{L}_{a}=-c_{2}^{2}\partial_{x}^{2}+\cos(\psi_{a}),$ (5.10)
with $\delta$-type boundary conditions on $(0,\infty)$.
Lemma 5.2.
Consider the operator $\mathcal{L}_{a}$ in (5.10) on the domain
$W_{Z}=\{g\in H^{2}(0,\infty):g^{\prime}(0+)=-Zg(0+)\}.$
Then, for $Z\in\mathbb{R}$ we have that $(\mathcal{L}_{a},W_{Z})$ represents a family of self-adjoint operators and that the Morse index of $\mathcal{L}_{a}$, $n(\mathcal{L}_{a})$, satisfies $n(\mathcal{L}_{a})\leqq 1$ for every $Z$.
Proof.
We will use the extension theory for symmetric operators (see Appendix D in [4]). In fact, by considering the minimal symmetric Schrödinger operator,
$\mathcal{M}_{0}\equiv\mathcal{L}_{a},\quad D(\mathcal{M}_{0})=C_{0}^{\infty}(0,\infty),$
we have that the closure of $\mathcal{M}_{0}$, denoted by $\mathcal{M}\equiv\mathcal{M}_{0}$, has deficiency indices $n_{\pm}(\mathcal{M})$, satisfying $n_{\pm}(\mathcal{M})=1$. By Proposition A.3 in Appendix we obtain that the family $(\mathcal{L}_{a},W_{Z})$ represents all the self-adjoint extensions of the closed symmetric operator $(\mathcal{M},D(\mathcal{M}))$. Next, we show that $\mathcal{M}\geqq 0$. Indeed, from (1.6) we obtain
$\mathcal{M}f(x)=-\frac{1}{\psi^{\prime}_{a}(x)}\frac{d}{dx}\Big{[}c_{2}^{2}(\psi^{\prime}_{a}(x))^{2}\frac{d}{dx}\Big{(}\frac{f(x)}{\psi^{\prime}_{a}(x)}\Big{)}\Big{]},\qquad x>0.$ (5.11)
We note that always we have $\psi^{\prime}_{a}\neq 0$ on $[0,\infty)$. Thus for $f\in C_{0}^{\infty}(0,\infty)$ we obtain
[…]
Due to the condition $f(0)=0$ the non-integral term vanishes, and we get $\mathcal{M}\geqq 0$. Then, by Proposition A.5 (see Appendix), we have all the self-adjoint extensions $(\mathcal{L}_{a},W_{Z})$ of $\mathcal{M}$ satisfy $n(\mathcal{L}_{a})\leqq 1$. This finishes the proof. ∎
Remark 5.3.
Notice that it is possible to have, in Lemma 5.2, that $n(\mathcal{L}_{a})=0$ for some $W_{Z}$-domain. Indeed, by considering $Z_{a}=-\psi^{\prime\prime}_{a}(0)/\psi^{\prime}_{a}(0)$ we have $\psi^{\prime}_{a}\in W_{Z_{a}}$ and $\mathcal{L}_{a}\psi^{\prime}_{a}=0$. Thus, as $\psi^{\prime}_{a}<0$ on $[0,\infty)$ follows from Sturm-Liouville oscillation theory on half-lines that zero is the smallest eigenvalue for $(\mathcal{L}_{a},W_{Z_{a}})$ (note that the essential spectrum for $\mathcal{L}_{a}$ is $\sigma_{\mathrm{ess}}(\mathcal{L}_{a})=[1,\infty)$).
By Theorem A.7 in Appendix we obtain that the Morse index of $(\mathcal{L},D_{Z,0})$ satisfies $n(\mathcal{L})\leqq 2$. Upon application of Lemmata 5.1 and 5.2 we obtain the following result.
Lemma 5.4.
Consider the pair $(\mathcal{L},D_{Z,0})$. Suppose that there are $(f_{i},g_{i})\in D_{Z,0}$ with $g_{i}(0)\neq 0$, $i=1,2$, such that $\mathcal{L}(f_{i},g_{i})^{\top}=\lambda_{i}(f_{i},g_{i})^{\top}$ with $\lambda_{i}<0$. Then, $\lambda_{1}=\lambda_{2}$ and $g_{1}=\alpha g_{2}$ on $[0,\infty)$.
Proof.
Suppose $\lambda_{1}<\lambda_{2}$. By Lemma 5.1 follows $\mathcal{L}_{a}g_{i}=\lambda_{i}g_{i}$, $g_{i}\in W_{Z}-\{0\}$. Therefore, $n(\mathcal{L}_{a})\geqq 2$, which is a contradiction by Lemma 5.2. Thus, $\lambda_{1}=\lambda_{2}$ and by Lemma 5.2 we need to have that $g_{1}$ and $g_{2}$ are linearly dependent. ∎
Remark 5.5.
We note from Lemma 5.1 that Lemma 5.4 implies $\mathcal{L}_{1}f_{i}=\lambda_{1}f_{i}$ with $f_{i}$ satisfying periodic boundary conditions; hence, it is possible for $\lambda_{1}$ to be a double eigenvalue.
Theorem 5.6.
Consider the pair $(\mathcal{L},D_{Z,0})$ determined by a priori $(\phi_{1},\psi_{a})$ positive single-lobe kink state, $a\in\mathbb{R}$. Then, the Morse index satisfies $n(\mathcal{L})\leqq 1$. Therefore, we have also that for the pair $(\mathcal{L}_{Z},D_{Z})$, $n(\mathcal{L}_{Z})\leqq 1$.
Proof.
By Theorem A.7 in Appendix, we known that $n(\mathcal{L})\leqq 2$. Suppose that $n(\mathcal{L})=2$ and consider $\lambda_{0},\lambda_{1}$ being the negative eigenvalues (counting multiplicity).
  1. (I)
    Suppose initially $\lambda_{0}<\lambda_{1}$. In the sequel we divide our analysis into several steps:
    1. (A)
      Perron-Frobenius property for $(\mathcal{L},D_{Z,0})$. For $\lambda_{0}=\text{inf}\;\sigma(\mathcal{L})<0$, let $(\xi_{\lambda_{0}},\zeta_{\lambda_{0}})$ be an associated eigenfunction to $\lambda_{0}$. Then, $\xi_{\lambda_{0}}$ and $\zeta_{\lambda_{0}}$ can be chosen as positive functions so that, in addition, $\xi_{\lambda_{0}}$ is even. Indeed, we consider the quadratic form $\mathcal{Q}_{Z}$ associated to operator $\mathcal{L}$ on $D_{Z,0}$, namely, $\mathcal{Q}_{Z}:D(\mathcal{Q}_{Z})\to\mathbb{R}$, with
      $\mathcal{Q}_{Z}(\xi,\zeta)=\int_{-L}^{L}(\xi^{\prime})^{2}+V_{\phi_{1}}\xi^{2}dx+\int_{0}^{\infty}(\zeta^{\prime})^{2}+W_{\psi_{a}}\zeta^{2}dx-Z|\zeta(0)|^{2},$ (5.13)
      $V_{\phi_{1}}=\cos(\phi_{1})$, $W_{\psi_{a}}=\cos(\psi_{a})$, and $D(\mathcal{Q}_{Z})$ defined by
      $D(\mathcal{Q}_{Z})=\{(\xi,\zeta)\in X^{1}(-L,L):\xi(L)=\xi(-L)=\zeta(0)\}.$ (5.14)
      1. (1)
        The profile $\zeta_{\lambda_{0}}$ is not identically zero. Indeed, suppose that $\zeta_{\lambda_{0}}\equiv 0$; then $\xi_{\lambda_{0}}$ satisfies
        $\left\{\begin{array}[]{ll}\mathcal{L}_{1}\xi_{\lambda_{0}}(x)=\lambda_{0}\xi_{\lambda_{0}}(x),\quad\quad x\in(-L,L),\\ \xi_{\lambda_{0}}(L)=\xi_{\lambda_{0}}(-L)=0,\\ \xi^{\prime}_{\lambda_{0}}(L)=\xi^{\prime}_{\lambda_{0}}(-L).\end{array}\right.$ (5.15)
        Next, from the Dirichlet condition (which implies that the eigenvalue $\lambda_{0}$ is simple and so $\xi_{\lambda_{0}}$ is odd or even) and from oscillations theorems of the Floquet theory (which imply that the number of zeros of $\xi_{\lambda_{0}}$ is even on $[-L,L)$), we need to have that $\xi_{\lambda_{0}}$ is odd. Then, by Sturm-Liouville theory there is an eigenvalue $\theta$ for $\mathcal{L}_{1}$, such that $\theta<\lambda_{0}$, with associated eigenfunction $\chi>0$ on $(-L,L)$, and $\chi(-L)=\chi(L)=0$. Now, let $\mathcal{Q}_{\mathrm{Dir}}$ be the quadratic form associated to $\mathcal{L}_{1}$ with Dirichlet domain, namely, $\mathcal{Q}_{\mathrm{Dir}}:H^{1}_{0}(-L,L)\to\mathbb{R}$ and
        $\mathcal{Q}_{\mathrm{Dir}}(f)=\int_{-L}^{L}(f^{\prime})^{2}+V_{\phi_{1}}f^{2}dx.$ (5.16)
        Then, $\mathcal{Q}_{\mathrm{Dir}}(\chi)=\mathcal{Q}_{Z}(\chi,0)\geqq\lambda_{0}\|\chi\|^{2}$ and so, $\theta\geqq\lambda_{0}$. This is a contradiction.
      2. (2)
        $\zeta_{\lambda_{0}}(0)\neq 0$: suppose $\zeta_{\lambda_{0}}(0)=0$ and we consider the odd-extension $\zeta_{\mathrm{odd}}$ for $\zeta_{\lambda_{0}}$, and the even-extension $\psi_{\mathrm{even}}$ of the kink-profile $\psi_{a}$ on all the line. Then, $\zeta_{\mathrm{odd}}\in H^{2}(\mathbb{R})$ and $\psi_{\mathrm{even}}\in H^{2}(\mathbb{R}-\{0\})\cap H^{1}(\mathbb{R})$. Next, we consider the unfold operator
        ~$\widetilde{\mathcal{L}}=-c_{2}^{2}\partial_{x}^{2}+\cos(\psi_{\mathrm{even}}),$ (5.17)
        on the $\delta$-interaction domain
        $D_{\delta,\gamma}=\{f\in H^{2}(\mathbb{R}-\{0\})\cap H^{1}(\mathbb{R}):f^{\prime}(0+)-f^{\prime}(0-)=\gamma f(0)\},$ (5.18)
        for any $\gamma\in\mathbb{R}$. Then, from the extension theory for symmetric operators we have that the family (~$(\widetilde{\mathcal{L}},D_{\delta,\gamma})_{\gamma\in\mathbb{R}}$ represents all the self-adjoint extensions of the symmetric operator $(\mathcal{N}_{0},D(\mathcal{N}_{0}))$ defined by
        $\mathcal{N}_{0}=\widetilde{\mathcal{L}},\quad D(\mathcal{N}_{0})=\{f\in H^{2}(\mathbb{R}):f(0)=0\}.$
        Moreover, the deficiency indices of $(\mathcal{N}_{0},D(\mathcal{N}_{0}))$ are given by $n_{\pm}(\mathcal{N}_{0})=1$ (see Albeverio et al. [4] and Proposition A.3 in Appendix below). Now, the even tail-profile $\psi_{\mathrm{even}}$ satisfies $\psi_{\mathrm{even}}^{\prime}(x)\neq 0$ for all $x\neq 0$, and so from the relation
        $\mathcal{N}_{0}f=-\frac{1}{\psi^{\prime}_{\mathrm{even}}}\frac{d}{dx}\Big{[}c_{2}^{2}(\psi^{\prime}_{\mathrm{even}})^{2}\frac{d}{dx}\Big{(}\frac{f}{\psi^{\prime}_{\mathrm{even}}}\Big{)}\Big{]},\quad x\in\mathbb{R}-\{0\},$ (5.19)
        we can see easily that $\langle\mathcal{N}_{0}f,f\rangle\geqq 0$ for all $f\in D(\mathcal{N}_{0})$. Then, from the extension theory (see Proposition A.5 in Appendix) we obtain that the Morse index for the family (~$(\widetilde{\mathcal{L}},D_{\delta,\gamma})$ satisfies $n(\widetilde{\mathcal{L}})\leqq 1$, for all $\gamma\in\mathbb{R}$. Therefore, since $\zeta_{\mathrm{odd}}\in D_{\delta,\gamma}$, for any $\gamma$, and ~$\widetilde{\mathcal{L}}\,\zeta_{\mathrm{odd}}=\lambda_{0}\zeta_{\mathrm{odd}}$ on $\mathbb{R}$, we have $n(\widetilde{\mathcal{L}})=1$ and $\lambda_{0}$ will be the smallest negative eigenvalue for $\widetilde{\mathcal{L}}$ on $\delta$-interactions domains. Then, by Theorem A.8 in Appendix (Perron Frobenius property for $\widetilde{\mathcal{L}}$), $\zeta_{\mathrm{odd}}$ needs to be positive which is a contradiction. Therefore, $\zeta_{\lambda_{0}}(0)\neq 0$.
      3. (3)
        $\zeta_{\lambda_{0}}$ and $\xi_{\lambda_{0}}$ are positive profiles: from $\zeta_{\lambda_{0}}(0)\neq 0$ and Lemma 5.1, the pair $(\zeta_{\lambda_{0}},\lambda_{0})$ satisfies the $(\delta BP)$-boundary problem in (5.6) and so by Lemma 5.2 we obtain $n(\mathcal{L}_{a})=1$. Therefore, by oscillation theory we can choose $\zeta_{\lambda_{0}}>0$ on $(0,\infty)$ (indeed, we can use the Perron-Frobenius property in Theorem A.8 for the unfold operator (~$(\widetilde{\mathcal{L}},D_{\delta,-Z})$ in (A.17)-(A.18) and also deduce the positive property of $\zeta_{\lambda_{0}}$). Next, we study $\xi_{\lambda_{0}}$. From Lemma 5.1, the pair $(\xi_{\lambda_{0}},\lambda_{0})$ satisfies the $(PBP)$-boundary problem in (5.6). We denote by $\eta_{0}$ the first eigenvalue (simple) for (PBP) in (5.6), then $\lambda_{0}\geqq\eta_{0}$. We are going to show that $\lambda_{0}=\eta_{0}$. Indeed, we consider the quadratic form, $\mathcal{Q}_{\mathrm{per}}$, associated to the (PBP)-problem in (5.6), namely, for $h\in H_{\mathrm{per}}^{1}(-L,L)$ (with $h(L)=h(-L),h^{\prime}(L)=h^{\prime}(-L)$) we have,
[…]
        Next, for $h$ fixed as above, define $\chi=\nu\zeta_{\lambda_{0}}$ with $\nu\in\mathbb{R}$ being chosen such that $\chi(0)=\nu\zeta_{\lambda_{0}}(0)=h(L)$. Then, $(h,\chi)\in D(\mathcal{Q}_{Z})$ in (5.14). Now, by using that $\mathcal{L}_{a}\zeta_{\lambda_{0}}=\lambda_{0}\zeta_{\lambda_{0}}$ on $(0,\infty)$, we obtain
        $\mathcal{Q}_{\mathrm{per}}(h)$ $=\mathcal{Q}_{Z}(h,\chi)-\int_{0}^{\infty}(\chi^{\prime})^{2}+W_{\psi_{a}}\chi^{2}dx+Z|\chi(0)|^{2}$ (5.21)
        $=\mathcal{Q}_{Z}(h,\chi)-\langle\mathcal{L}_{a}\chi,\chi\rangle=\mathcal{Q}_{Z}(h,\chi)-\lambda_{0}\|\chi\|^{2}$
        $\geqq\lambda_{0}[\|h\|^{2}+\|\chi\|^{2}]-\lambda_{0}\|\chi\|^{2}=\lambda_{0}\|h\|^{2}.$
        Then, $\eta_{0}\geqq\lambda_{0}$ and so $\eta_{0}=\lambda_{0}$. Therefore, from Floquet theory follows that we can choose $\xi_{\lambda_{0}}$ positive on $[-L,L]$. Even more, since $V_{\phi_{1}}$ is even, we obtain that $\xi_{\lambda_{0}}$ is even.
    2. (B)
      Let $(f_{{1}},g_{1})\in D_{Z,0}$ be an associated eigenfunction for $\lambda_{{1}}$. We shall prove that $g_{1}(0+)\neq 0$.
      1. (1)
        Suppose that $g_{{1}}\equiv 0$: then $f_{{1}}(-L)=f_{{1}}(L)=0$, $f^{\prime}_{{1}}(-L)=f^{\prime}_{{1}}(L)$, and so $f_{1}$ will be odd (see step 1 in item A) above). Now, our single-lobe profile $\phi_{1}$ satisfies $\mathcal{L}_{1}\phi^{\prime}_{1}=0$, $\phi^{\prime}$ is odd, $\phi^{\prime}(x)>0$ for $x\in[-L,0)$, thus since $\lambda_{1}<0$ we obtain from the Sturm oscillation theorem that there is $r\in(-L,0)$ such that $\phi^{\prime}(r)=0$, it which is a contradiction. Then, $g_{1}$ is non-trivial.
      2. 2)
        Now suppose $g_{1}(0+)=0$: we consider the odd-extension $g_{1,\mathrm{odd}}\in H^{2}(\mathbb{R})$ of $g_{1}$ on all the line and the extension operator $\widetilde{\mathcal{L}}$ in (5.17) on the $\delta$-interaction domains $D_{\delta,\gamma}$ in (5.18). Then, $g_{1,\mathrm{odd}}\in D_{\delta,\gamma}$ for any $\gamma$ and by Perron-Frobenius property (Theorem A.8) we will obtain $n(\widetilde{\mathcal{L}})\geqq 2$, it which is a contradiction because $n(\widetilde{\mathcal{L}})\leqq 1$ for all $\gamma$ (see item (A)(2) above).
    3. (C)
      By items (A), (B) and Lemma 5.4 we have that $\lambda_{1}=\lambda_{0}$, which is a contradiction.
  2. (II)
    Suppose $\lambda_{0}$ is a double eigenvalue: $\lambda_{0}=\lambda_{1}$. Then by Perron-Frobenius property in item (I)(A) there holds $\xi_{\lambda_{0}},\zeta_{\lambda_{0}}>0$. Now, for $(f_{1},g_{1})\in D_{Z,0}$ an associated eigenfunction for $\lambda_{1}$, we have $(f_{1},g_{1})\perp(\xi_{\lambda_{0}},\zeta_{\lambda_{0}})$. Next, by item (I)(B), Lemma 5.1 and $\lambda_{0}=\eta_{0}$, being $\eta_{0}$ the first eigenvalue (simple) for (PBP) in (5.6), we have that $f_{1}=r\xi_{\lambda_{0}}$ and $g_{1}=s\zeta_{\lambda_{0}}$ with $r=s\neq 0$ by continuity at zero. Hence, we arrive to a contradiction from the orthogonality property of the eigenfunctions. This finishes the proof of the Theorem.
Next, we will give sufficient conditions for obtaining $n(\mathcal{L}_{Z})=1$, with $\mathcal{L}_{Z}$ defined in (5.1) and $(\phi_{1},\phi_{2})\equiv(\phi_{1,k},\phi_{a})$, with the single-lobe state $\phi_{1,k}$ in (4.7) or (4.8). Initially, we consider the mapping $\mathcal{T}(\theta)=\theta\cos\theta-\sin\theta$ for $\theta\in[0,2\pi]$. Then $\mathcal{T}(\theta)\leqq 0$ for $\theta\in[0,\theta_{0}]$ where $\theta_{0}\approx 4.4934$ is the unique zero of $\mathcal{T}$ on the interval $(0,2\pi)$. We note that $\theta_{0}\in(\pi,\frac{3\pi}{2})$ (see Figure 19).
Figure 19: Graph of $\mathcal{T}(\theta)=\theta\cos\theta-\sin\theta$, $\theta\in[0,2\pi]$.
Thus, with the notations above we get the following results.
Lemma 5.7.
Consider the single-lobe state $\phi_{1,k}$ determined by $k\in(0,1)$ in (4.7) or (4.8). Then, for $k$ such that
$k^{2}\leqq\frac{1+\cos(\theta_{0})}{2}\approx 0.3914,$
we have that $\phi_{1,k}(0)\in(\pi,\theta_{0}]$ and, therefore,
$\mathcal{T}(\phi_{1,k}(x))=\phi_{1,k}(x)\;\cos(\phi_{1,k}(x))-\sin(\phi_{1,k}(x))\leqq 0,\qquad\text{for all}\;\;x\in[-L,L].$ (5.22)
Moreover, for the center solution $\phi_{1}(x)\equiv\pi$, relation in (5.22) is trivial.
Proof.
By (4.7) or (4.8) and $\operatorname{sn}^{2}(K(k);k)=1$, for all $k$, we have $\phi_{1,k}(0)=2\pi-\arccos(-1+2k^{2})\leqq\theta_{0}$ if and only if
$\cos(2\pi-\arccos(2k^{2}-1))=2k^{2}-1\leqq\cos\theta_{0}.$
We recall that $\phi_{1,k}(0)\in(\pi,2\pi)$. Lastly, since $\phi_{1,k}(x)\in(0,\phi_{1,k}(0)]$ for all $x\in[-L,L]$, we get immediately (5.22). This finishes the proof. ∎
Theorem 5.8.
Consider the single-lobe state $\phi_{1,k}$ in (4.7) or (4.8) determined by $k\in(0,1)$ such that
$k^{2}\leqq\frac{1+\cos(\theta_{0})}{2}\approx 0.3914,$
and such that for this value of $k$ we have the existence of a single-lobe kink state $(\phi_{1,k},\phi_{2,a})\in D_{Z}$, with $a=a(k)$. Then for $(\mathcal{L}_{Z},D_{Z})$ defined in (5.1) with $(\phi_{1},\phi_{2})\equiv(\phi_{1,k},\phi_{2,a})$, we have that the Morse index satisfies $n(\mathcal{L}_{Z})=1$. In the case of the degenerate single-lobe kink state $(\pi,\phi_{2,0})\in D_{Z}$ with $Z=\frac{2}{\pi c_{2}}$ we also have that $n(\mathcal{L}_{Z})=1$.
Proof.
For the single-lobe kink state $(\phi_{1,k},\phi_{2,a})\in D_{Z}$, we have for $x>L$ that $\phi_{2,a}(x)\leqq\phi_{2,a}(L)=\phi_{1,k}(L)<\theta_{0}$. Then, for all $x>L$ we obtain $\mathcal{T}(\phi_{2,a}(x))=-\sin(\phi_{2,a}(x))+\phi_{2,a}(x)\cos(\phi_{2,a}(x))<0$. Therefore, by Lemma 5.7 we obtain
[…]
Then from minimax principle we arrive at $n(\mathcal{L}_{Z})\geqq 1$ (similarly for the degenerate case $(\pi,\phi_{2,0})$). Thus, by Theorem 5.6 we finish the proof. ∎
Remark 5.9.
  1. (1)
    Notice that by Theorem 5.8 and item (A) in the proof of Theorem 5.6, we obtain that the unique negative eigenvalue for $(\mathcal{L}_{Z},D_{Z})$ coincides with the smallest negative eigenvalue for the eigenvalues problems (PBP) and ($\delta$BP) in (5.6).
  2. (2)
    For $k^{2}>\frac{1+\cos(\theta_{0})}{2}$ our study does not provide an accurate picture of stability properties of the possible profiles $(\phi_{1,k},\phi_{2,a})$.
In what follows we exactly determine which single-lobe kink state $(\phi_{1,k},\phi_{2,a})\in D_{Z}$, characterized by Propositions 4.7 and 4.9, will induce that the Morse index of $\mathcal{L}_{Z}$ is equal to one. As we will see the value of $L$, $k$, $c_{i}$ and the sign of $Z$ play a fundamental role.
Theorem 5.10.
Consider the single-lobe state $\phi_{1,k}$ in (4.7) such that the single-lobe kink state $(\phi_{1,k},\phi_{2,a})\in D_{Z}$, $a=a(k)<0$ (see formula (4.11)), it is determined by Proposition 4.7. Then the following conditions on $L$, $c_{i}$, $Z$, and with $k$ satisfying
$k^{2}\leqq\frac{1+\cos(\theta_{0})}{2}\equiv k^{2}_{\ell}\approx 0.3914,$
imply that the operator $\mathcal{L}_{Z}$ in (5.1) associated to $(\phi_{1,k},\phi_{2,a(k)})$ satisfies $n(\mathcal{L}_{Z})=1$. To state the conditions we denote $\mathcal{I}=K(k_{\ell})\approx 1.77160$ and for $L,c_{1}$ fixed but arbitrary we obtain:
  1. (1)
    Let $\frac{L}{c_{1}}>\frac{\pi}{2}$ and $k_{0}\in(0,1)$ such that $K(k_{0})=\frac{L}{c_{1}}$. Then for $\frac{L}{c_{1}}<\mathcal{I}$ we consider the family of solutions
    $k\in(k_{0},k_{\ell})\to(\phi_{1,k},\phi_{2,a}),$
    with $a=a(k)$ determined by formula in (4.11). Then for $H(k)$ defined in (4.17) we have:
    1. (i)
      For $Z>0$ with $Z\in(H(k_{\ell}),H(k_{0}))$, $H(k_{0})=\frac{4}{\pi c_{1}}\big{[}\frac{c_{1}}{2c_{2}}-k_{0}\big{]}$, and with $c_{2}$ satisfying $\frac{c_{1}}{2c_{2}}\geqq\tanh(\frac{L}{c_{1}})$ we have $n(\mathcal{L}_{Z})=1$.
    2. (ii)
      let $Z>0$ and $g(k)=k\operatorname{sn}(L/c_{1};k)$, $k\in(k_{0},1)$, suppose that $c_{2}$ satisfies $g(k_{\ell})<\frac{c_{1}}{2c_{2}}<tanh(\frac{L}{c_{1}})$. Then for $Z\in(H(k_{\ell}),H(k_{0}))$ we have $n(\mathcal{L}_{Z})=1$.
    3. (iii)
      Let $Z<0$ and suppose that $c_{2}$ satisfies $\frac{c_{1}}{2c_{2}}<k_{0}$. Then for $Z\in(H(k_{0}),H(k_{\ell}))$ we have $n(\mathcal{L}_{Z})=1$.
    4. (iv)
      Let $Z<0$ and suppose that $c_{2}$ satisfies $\frac{c_{1}}{2c_{2}}=k_{0}$. Then for $Z\in(H(k_{\ell}),0)$ we have $n(\mathcal{L}_{Z})=1$.
    5. (v)
      Let $Z<0$, $g(k)=k\operatorname{sn}(L/c_{1};k)$, $k\in(k_{0},1)$ and suppose that $c_{2}$ satisfies $g(k_{\ell})>\frac{c_{1}}{2c_{2}}>k_{0}$. Then for $Z\in(H(k_{\ell}),0)$ we have $n(\mathcal{L}_{Z})=1$.
    6. (vi)
      Let $Z=0$ and suppose that $c_{2}$ satisfies $\frac{c_{1}}{2c_{2}}\in(k_{0},k_{\ell}\operatorname{sn}(L/c_{1};k_{\ell}))$. Then $n(\mathcal{L}_{0})=1$.
  2. 2)
    Let $\frac{L}{c_{1}}\leqq\frac{\pi}{2}$ and consider the family of solutions
    $k\in(0,k_{\ell})\to(\phi_{1,k},\phi_{2,a}),$
    with $a=a(k)$ determined by formula in (4.11). Then, for $H(k)$ defined in (4.17) we have the following:
    1. (i)
      For $Z>0$ we have two cases:
      1. (a)
        Suppose $c_{2}$ satisfies $\frac{c_{1}}{2c_{2}}\geqq\tanh(\frac{L}{c_{1}})$. Then for $Z\in(H(k_{\ell}),\frac{2}{\pi c_{2}})$ we have $n(\mathcal{L}_{Z})=1$.
      2. (b)
        Suppose $c_{2}$ satisfies $\frac{c_{1}}{2c_{2}}<\tanh(\frac{L}{c_{1}})$ and let $\beta\in(0,1)$ be such that $g(\beta)=\frac{c_{1}}{2c_{2}}$. Then if $k_{\ell}<\beta$ we have for all $Z\in(H(k_{\ell}),\frac{2}{\pi c_{2}})$ that $n(\mathcal{L}_{Z})=1$.
    2. (ii)
      Let $Z<0$ and suppose that $c_{2}$ satisfies $k_{\ell}\operatorname{sn}(L/c_{1};k_{\ell})>\frac{c_{1}}{2c_{2}}$. Then for $Z\in(H(k_{\ell}),0)$ we have $n(\mathcal{L}_{Z})=1$.
    3. (iii)
      Let $Z=0$ and suppose that $c_{2}$ satisfies $\frac{c_{1}}{2c_{2}}\in(0,k_{\ell}\operatorname{sn}(L/c_{1};k_{\ell}))$. Then $n(\mathcal{L}_{0})=1$.
Proof.
By Proposition 4.7 and its proof we have the following cases:
  1. (1)
    Suppose $\frac{L}{c_{1}}>\frac{\pi}{2}$:
    1. (i)
      For $Z>0$ and $Z\in(H(k_{\ell}),H(k_{0}))$ it follows from the strictly-decreasing property of the mapping $k\in(k_{0},k_{\ell})\to H(k)$ and the condition on $c_{2}$ that there is $k_{Z}\in(k_{0},k_{\ell})$ such that $H(k_{Z})=Z$. Therefore, $\Phi_{k_{Z}}=(\phi_{1,k_{Z}},\phi_{2,a})\in D_{Z}$, $a=a(k_{Z})$ and the operator $\mathcal{L}_{Z}$ associated to $\Phi_{k_{Z}}$ satisfies $n(\mathcal{L}_{Z})=1$.
    2. ii)
      Let $Z>0$. By the conditions on $c_{2}$ and $g(k_{0})=k_{0}<g(k_{\ell})$ we get $\frac{c_{1}}{2c_{2}}\in(k_{0},\tanh(\frac{L}{c_{1}}))$ and $k_{\ell}\in(k_{0},\gamma)$ with $g(\gamma)=\frac{c_{1}}{2c_{2}}$ and $H(\gamma)=0$. Therefore, there is $k_{Z}\in(k_{0},k_{\ell})$ such that $H(k_{Z})=Z$ and the profile $\Phi_{k_{Z}}=(\phi_{1,k_{Z}},\phi_{2,a})\in D_{Z}$, $a=a(k_{Z})$. Hence, $\mathcal{L}_{Z}$ associated to $\Phi_{k_{Z}}$ satisfies $n(\mathcal{L}_{Z})=1$.
    3. (iii)
      Let $Z<0$. By the condition on $c_{2}$ and the strictly-increasing property of the mapping $k\in(k_{0},k_{\ell})\to H(k)$ there is $k_{Z}\in(k_{0},k_{\ell})$ such that $H(k_{Z})=Z$. Thus, we conclude that $n(\mathcal{L}_{Z})=1$.
    4. (iv)
      Let $Z<0$. By the condition on $c_{2}$ we obtain $H(k_{0})=0$. Thus, as $H(k_{\ell})\geqq m_{0}=\min_{k\in(k_{0},1)}H(k)$ we have for $Z\in(H(k_{\ell}),0)$ that there is $k_{Z}\in(k_{0},k_{\ell})$ such that $H(k_{Z})=Z$. Then, $n(\mathcal{L}_{Z})=1$.
    5. (v)
      Let $Z<0$. By the condition on $c_{2}$ and $k\operatorname{sn}(L/c_{1};k)<\tanh(\frac{L}{c_{1}})$, we obtain $\frac{c_{1}}{2c_{2}}\in\big{(}k_{0},\tanh(\frac{L}{c_{1}})\big{)}$. Next, let $\beta\in(k_{0},k_{\ell})$ be such that $g(\beta)=\frac{c_{1}}{2c_{2}}$ and, thus, $H(\beta)=0$. Then for $Z\in(H(k_{\ell}),0)$ there is $k_{Z}\in(\beta,k_{\ell})$ such that $H(k_{Z})=Z$. We conclude that $n(\mathcal{L}_{Z})=1$.
    6. (vi)
      Let $Z=0$. By the condition on $c_{2}$ there is $k\in(k_{0},k_{\ell})$ such that $k\operatorname{sn}(L/c_{1};k)=\frac{c_{1}}{2c_{2}}$. Then $H(k)=0$ and the operator $\mathcal{L}_{0}$ associated to $(\phi_{1,k},\phi_{2,a(k)})$ satisfies $n(\mathcal{L}_{0})=1$.
  2. 2)
    Now suppose that $\frac{L}{c_{1}}\leqq\frac{\pi}{2}$:
    1. (i)
      For $Z>0$ we have the following cases:
      1. (a)
        Suppose that $c_{2}$ satisfies $\frac{c_{1}}{2c_{2}}\geqq\tanh(\frac{L}{c_{1}})$. Then $H(k)>0$ for all $k\in(0,1)$ with $H(0)=\frac{2}{\pi c_{2}}$, $H(1)=0$ and $H^{\prime}(k)<0$. Thus for $Z\in(H(k_{\ell}),\frac{2}{\pi c_{2}})$ there is $k_{Z}\in(0,k_{\ell})$ such that $H(k_{Z})=Z$. We conclude that $n(\mathcal{L}_{Z})=1$.
      2. (b)
        Suppose that $c_{2}$ satisfies $\frac{c_{1}}{2c_{2}}<\tanh(\frac{L}{c_{1}})$ and let $\beta\in(0,1)$ be such that $g(\beta)=\frac{c_{1}}{2c_{2}}$. Then $H(\beta)=0$ and $H(k)>0$ for all $k\in(0,\beta)$ with $H(0)=\frac{2}{\pi c_{2}}$ and $H^{\prime}(k)<0$. Thus, if $k_{\ell}<\beta$ we have for all $Z\in(H(k_{\ell}),\frac{2}{\pi c_{2}})$ that $n(\mathcal{L}_{Z})=1$.
    2. (ii)
      For $Z<0$ and $c_{2}$ satisfying $k_{\ell}\operatorname{sn}(L/c_{1};k_{\ell})>\frac{c_{1}}{2c_{2}}$, let $\beta\in(0,1)$ be such that $g(\beta)=\frac{c_{1}}{2c_{2}}$. Then, $H(\beta)=H(1)=0$, $H(k)<0$ for $k\in(\beta,1)$ and for $Z\in(H(k_{\ell}),0)$ we have $n(\mathcal{L}_{Z})=1$.
    3. (iii)
      For $Z=0$ and $c_{2}$ satisfying $\frac{c_{1}}{2c_{2}}\in(0,k_{\ell}\operatorname{sn}(L/c_{1};k_{\ell}))$, we obtain $k\in(0,k_{\ell})$ such that $k\operatorname{sn}(L/c_{1};k)=\frac{c_{1}}{2c_{2}}$ (we recall that $g(k)=k\operatorname{sn}(L/c_{1};k)$ is strictly-increasing) and hence $H(k)=0$. Then $n(\mathcal{L}_{0})=1$.
This finishes the examination of all the cases and the proof. ∎
Remark 5.11.
For $\frac{L}{c_{1}}>\frac{\pi}{2}$ and $k_{0}\in(0,1)$ such that $K(k_{0})=\frac{L}{c_{1}}$, we obtain that for $\frac{L}{c_{1}}\geqq I=K(k_{\ell})$ that $k_{0}>k_{\ell}$ and for the family $k\in(k_{0},1)\to(\phi_{1,k},\phi_{2,a})$ we can not apply Theorem 5.8.
Theorem 5.12.
Let $L,c_{1}$ be arbitrary but fixed with $\frac{L}{c_{1}}>\frac{\pi}{2}$ and $k_{0}$ such that $K(k_{0})=\frac{L}{c_{1}}$. Let $c_{2}$ satisfy
$\frac{c_{1}}{2k_{0}}<c_{2}.$
Then, for $k\in(0,k_{0})$, by Proposition 4.9 there exists a shift-value $a=a(k)>0$ (see Formula (4.11)), such that the single-lobe kink state satisfies $(\phi_{1,k},\phi_{2,a})\in D_{0}$, with $\phi_{1,k}$ in (4.8). Then the following restrictions on $L$, $c_{1}$ and with $k$ satisfying
$k^{2}\leqq\frac{1+\cos(\theta_{0})}{2}\equiv k^{2}_{\ell}\approx 0.3914,$
guarantee that the operator $\mathcal{L}_{0}$ in (5.1) associated to $(\phi_{1,k},\phi_{2,a(k)})$ satisfies $n(\mathcal{L}_{0})=1$. In order to state these conditions, let $\mathcal{I}=K(k_{\ell})\approx 1.77160$; then we have:
  1. (1)
    For $\frac{L}{c_{1}}<\mathcal{I}$, the self-adjoint operator $\mathcal{L}_{0}$ in (5.1) determined by the family of solutions $k\in(0,k_{0})\to(\phi_{1,k},\phi_{2,a})$ satisfies $n(\mathcal{L}_{0})=1$.
  2. (2)
    For $\frac{L}{c_{1}}>\mathcal{I}$, the self-adjoint operator $\mathcal{L}_{0}$ in (5.1) determined by the family of solutions $k\in(0,k_{\ell})\to(\phi_{1,k},\phi_{2,a})$ satisfies $n(\mathcal{L}_{0})=1$.
Proof.
The proof is an immediate consequence of Theorem 5.8. ∎
5.2 Trivial kernel for $\mathcal{L}_{Z}$
The purpose of this subsection is to provide sufficient conditions for $\mathcal{L}\equiv(\mathcal{L}_{1},\mathcal{L}_{a})$ to have a trivial kernel, with $\mathcal{L}_{1},\mathcal{L}_{a}$ defined in (5.3) and $(\phi_{1},\psi_{a})$ being an a priori positive single-lobe kink with
$\psi_{a}(x)=4\arctan\Big{[}e^{-\frac{1}{c_{2}}(x+a)}\Big{]},\qquad x>0.$
Theorem 5.13.
Let $Z\in(-\infty,\frac{2}{\pi c_{2}})$ be an admissible strength value to obtain a single-lobe kink state $(\phi_{1},\psi_{a})\in D_{Z,0}$ and define the quantity
$\alpha_{a}\equiv\frac{\psi^{\prime\prime}_{a}(0)}{\psi^{\prime}_{a}(0)}=-\frac{1}{c_{2}}\tanh\Big{(}\frac{a}{c_{2}}\Big{)}.$
Then, the kernel associated to $\mathcal{L}=(\mathcal{L}_{1},\mathcal{L}_{a})$ on $D_{Z,0}$ is trivial in the following cases:
  1. (1)
    for $\alpha_{a}\neq-Z$, or
  2. (2)
    for $\alpha_{a}=-Z$ in the case of admissible $Z$ satisfying $Z\leqq 0$.
In particular,
  1. (i)
    for the case $Z\neq 0$ and $a<0$, the kernel of $\mathcal{L}$ is trivial;
  2. (ii)
    for the case $Z<0$ and $a>0$, the kernel of $\mathcal{L}$ is trivial; and,
  3. (iii)
    for the case $Z=0$, the kernel of $\mathcal{L}$ is trivial for any $a\in\mathbb{R}-\{0\}$.
In the case of a degenerate single-lobe kink state, $(\pi,\phi_{2,0})\in D_{Z}$ with $Z=\frac{2}{\pi c_{2}}$, we also have that $\ker(\mathcal{L}_{Z})$ is trivial.
Next, as an application of Theorem 5.13 and from Theorems 5.10 and 5.12, we establish our main result about the Morse and nullity indices for $\mathcal{L}_{Z}$.
Theorem 5.14.
Let $Z\in(-\infty,\frac{2}{\pi c_{2}})$ be an admissible strength value to obtain a single-lobe kink state $(\phi_{1,k},\psi_{2,a})\in D_{Z}$, given by Propositions 4.7 and 4.9. Then, under the restrictions on $L,c_{1},c_{2},k$ in Theorem 5.10 and 5.12, we obtain $n(\mathcal{L}_{Z})=1$ and $\ker(\mathcal{L}_{Z})=\{\bf{0}\}$. In the case of a degenerate single-lobe kink state $(\pi,\phi_{2,0})\in D_{Z}$ with $Z=\frac{2}{\pi c_{2}}$, we also have that $\ker(\mathcal{L}_{Z})$ is trivial and $n(\mathcal{L}_{Z})=1$.
Proof.
The proof is immediate from Theorems 5.10 and 5.12, and items $(i)$ and $(iii)$ in Theorem 5.13. ∎
Proof of Theorem 5.13.
The proof follows a similar analysis as in Angulo [6, 7]. By convenience of the reader we provide the main point in the analysis. Thus, let $(f,h)\in D_{Z,0}$ such that $\mathcal{L}(f,h)^{\top}={\bf{0}}$. Then $\mathcal{L}_{a}h(x)=0$ for $x>0$. Next, from classical Sturm-Liouville theory on half-line (see [19]) and $\mathcal{L}_{a}\psi^{\prime}_{a}=0$, we have $h=c\psi^{\prime}_{a}$ on $(0,\infty)$ for $c\in\mathbb{R}$. Now, we consider the following cases:
  1. (I)
    Suppose that $c=0$. This implies that $h\equiv 0$ and hence $f$ satisfies $\mathcal{L}_{1}f=0$ with Dirichlet-periodic boundary conditions on $[-L,L]$, $f(L)=f(-L)=0$ and $f^{\prime}(L)=f^{\prime}(-L)$. Suppose that $f\neq 0$:
    1. (a)
      From Sturm-Liouville theory for Dirichlet conditions we have that $f$ is even or odd (as zero is a simple eigenvalue and the potential $V_{\phi_{1}}$ is even).
    2. (b)
      Suppose that $f$ is odd: then, since $\mathcal{L}_{1}\phi^{\prime}_{1}=0$ and $\phi^{\prime}_{1}$ is odd, there is $\gamma\in\mathbb{R}$ such that $f=\gamma\phi^{\prime}_{1}$ (classical ODE analysis for second-order differential equations). Therefore, $0=f(L)=\gamma\phi^{\prime}_{1}(L)$ implies $\gamma=0$ (because $\phi^{\prime}_{1}(L)\neq 0$). Then, $f\equiv 0$.
    3. (c)
      Suppose that $f$ is even: then from oscillation results of the Floquet theory, we need to have that zero it is not the first eigenvalue for $\mathcal{L}_{1}$. Therefore, $f$ needs to change of sign and the number of zeros of $f$ on $[-L,L)$ is even. Hence $f$ has an odd number of zeros on $(-L,L)$ (as $f(-L)=0$). Then, $f(0)=0$ and $f^{\prime}(0)=0$. Hence, $f\equiv 0$.
    Thus, from items $(a),(b)$ and $(c)$ we obtain $f\equiv 0$ and $\ker(\mathcal{L})=\{\bf{0}\}$.
  2. (II)
    Suppose that $\alpha_{a}\neq-Z$. Then $c=0$ and by the former item (I) it follows that $\ker\mathcal{L})=\{\bf{0}\}$. Indeed, if $c\neq 0$ then $h(0)\neq 0$ and from Lemma 5.1 we get the eigenvalue problems in (5.6). In particular, $h$ and $Z$ satisfy $h^{\prime}(0)=-Zh(0)$ and hence $Z+\frac{\psi^{\prime\prime}_{a}(0)}{\psi^{\prime}_{a}(0)}=0$, which is a contradiction.
  3. (III)
    Suppose that $\alpha_{a}=-Z$ with $Z\leqq 0$. Then $c=0$ and so by the previous item (I) it follows that $\ker(\mathcal{L})=\{\bf{0}\}$. Suppose that $c\neq 0$; without loss of generality we can assume that $c<0$. Hence, $h(x)>0$ on $[0,\infty)$ and by the splitting eigenvalue result in Lemma 5.1 we obtain that $(f,h)$ satisfies the separated eigenvalues problems (PBP)-($\delta$BP) in (5.6) with $\lambda=0$. The $\delta$-interaction condition (also by hypotheses) implies
    $Z=-\frac{\psi^{\prime\prime}_{a}(0)}{\psi^{\prime}_{a}(0)}.$ (5.23)
    Next, we denote by $\{\eta_{n}\}$, $n\in\mathbb{N}_{0}$, the set of eigenvalues for the periodic-problem associated to $\mathcal{L}_{1}$ on $[-L,L]$, and by $\{\mu_{n}\}$, $n\in\mathbb{N}_{0}$, the set of eigenvalues for the Dirichlet-problem associated to $\mathcal{L}_{1}$ on $[-L,L]$. Then, by classical Sturm-Liouville theory (see Theorem 4.8.1 in [52]) we have that $\eta_{0}$ is simple and, in particular, we have the following distribution of eigenvalues,
    $\eta_{0}<\mu_{0}<\eta_{1}\leqq\mu_{1}\leqq\eta_{2}<\mu_{2}<\eta_{3}.$ (5.24)
    Now we prove that $\eta_{1}=0$ in (5.24) and that it is simple. Indeed, we examine all the cases:
    1. (a)
      Suppose that $0>\mu_{1}$. We know that $\mathcal{L}_{1}\phi_{1}^{\prime}=0$, $\phi^{\prime}_{1}$ is odd and strictly decreasing on $[-L,0]$. Now, since the eigenfunction $\chi$ associated to $\mu_{1}$ is odd and $\chi(-L)=\chi(0)=0$, it follows from the Sturm Comparison Theorem (cf. [51, 52]) that $\phi^{\prime}_{1}$ needs to have one zero on $(-L,0)$. But this is impossible and we have a contradiction.
    2. (b)
      Suppose that $\mu_{1}=0$. Then we have the existence of one odd-eigenfunction $\chi$ for $\mu_{1}$ ($\mathcal{L}_{1}\chi=0$, $\chi(-L)=\chi(L)=0$). Hence, by classical arguments from the theory of second-order differential equations, we get $\chi=\gamma\phi^{\prime}_{1}$ and $0=\chi(L)=\gamma\phi^{\prime}_{1}(L)$. Hence, $\gamma=0$ and $\chi\equiv 0$. Again, this is not possible and we reach a contradiction.
    3. (c)
      From the previous items (a) and (b) it follows that $\mu_{1}>0$ and, by (5.24) and Lemma 5.1 ($\mathcal{L}_{1}f=0$), we need to have $\eta_{1}=0$ and that the zero is simple. Moreover, the eigenfunction $f$ is even (for $f$ being odd, we have $f(L)=-f(-L)=-f(L)$ and so $h(0)=f(L)=0$), which is a contradiction.
    Next, by Floquet theory, $f$ has exactly two different zeros $-\beta,\beta$ ($\beta>0$) on $(-L,L)$. Therefore, $f(0)<0$, $f(\pm L)=h(0)>0$). Now, we consider the Wronskian function (constant) of $f$ and $\phi^{\prime}_{1}$ (we recall $\mathcal{L}_{1}f=\mathcal{L}_{1}\phi^{\prime}_{1}=0$),
    $W(x)=f(x)\phi^{\prime\prime}_{1}(x)-f^{\prime}(x)\phi^{\prime}_{1}(x)\equiv C,\quad\text{for all}\;\;x\in[-L,L].$
    Then, $C=f(0)\phi^{\prime\prime}_{1}(0)>0$. Now, by (5.23), $f^{\prime}(L)=0$, $f(L)=h(0)$, and from (1.6) we obtain
    $C=f(L)\phi^{\prime\prime}_{1}(L)=h(0)\psi^{\prime\prime}_{a}(0)=-cZ[\psi^{\prime}_{a}(0)]^{2}\leqq 0,$ (5.25)
    with $Z\leqq 0$. Therefore we get a contradiction by assuming that $c\neq 0$.
The statements $(i)-(iii)$ are immediate from the explicit formula for $\alpha_{a}$. This finishes the proof. ∎
Next, we establish our main result about the instability of a family of single-lobe kink states for the sine-Gordon model on a tadpole graph.
Theorem 5.15.
Under the restrictions on $Z,L,c_{1},c_{2}$ and $k$ given by Propositions 4.7 and 4.9 and by Theorems 5.10, 5.12 and 5.14, every member of the smooth family of single-lobe kink states, $k\to(\phi_{1,k},\phi_{2,a(k)})\in D_{Z}$, is spectrally and nonlinearly unstable for the sine-Gordon model (1.4) on a tadpole graph. The degenerate single-lobe kink state $(\pi,\phi_{2,0})\in D_{Z}$, with $Z=\frac{2}{\pi c_{2}}$ and
$\phi_{2,0}(x)=4\arctan\big{[}e^{-\frac{1}{c_{2}}(x-L)}\big{]},\qquad x\in[L,\infty),$
is also spectrally and nonlinearly unstable for the sine-Gordon model (1.4) on a tadpole graph.
Proof.
The proof of the linear instability property of the single-lobe kink states $(\phi_{1,k},\phi_{2,a(k)})$ is immediate from Theorem 5.14 and the instability criterion in Theorem 3.2. Now, since the mapping data-solution for the sine-Gordon model on the energy space $\mathcal{E}(\mathcal{G})\times L^{2}(\mathcal{G})$ is at least of class $C^{2}$ (indeed, it is smooth) by Theorem 2.1, it follows that the linear instability property of $(\phi_{1,k},\phi_{2,a(k)})$ is in fact of nonlinear type in the $\mathcal{E}(\mathcal{G})\times L^{2}(\mathcal{G})$-norm. The reader is referred to [10, 13, 30] for further information. This finishes the proof. ∎
Acknowledgements
J. Angulo was supported in part by Universal/CNPq project and by CNPq/Brazil Grant. The work of R. G. Plaza was partially supported by DGAPA-UNAM, program PAPIIT, grant IN-104922. R. G. Plaza is grateful to the IME-USP for their hospitality and to the FAPESP, Brazil (processo 2024/15816-0) for their financial support during a research stay in November 2024 when this work was carried out.
Appendix A Appendix
In this section we formulate some tools from the extension theory of symmetric operators by Krein and von Neumann which are suitable for our needs (see, for instance, Naimark [37, 38] and Reed and Simon [44] for further information). In particular, we establish for (~$(\widetilde{\mathcal{L}},D_{\delta,\gamma})_{\gamma\in\mathbb{R}}$ in (5.17) and (5.18), a Perron-Frobenius result to be used in the accurate Morse index estimate $n(\mathcal{L}_{Z})\leqq 1$ for $(\mathcal{L}_{Z},D_{Z})$ in (5.1) and (1.7). We show, for completeness in the exposition, the general estimate $n(\mathcal{L}_{Z})\leqq 2$ using extension theory.
A.1 Extension theory for the Laplacian operator on a tadpole graph
The following three results from the extension theory of symmetric operators are classical and can be found in [44].
Theorem A.1 (von-Neumann decomposition).
Let $A$ be a closed, symmetric operator, then
$D(A^{*})=D(A)\oplus\mathcal{N}_{-i}\oplus\mathcal{N}_{+i}.$ (A.1)
with $\mathcal{N}_{\pm i}=\ker(A^{*}\mp iI)$. Therefore, for $u\in D(A^{*})$ and $u=x+y+z\in D(A)\oplus\mathcal{N}_{-i}\oplus\mathcal{N}_{+i}$,
$A^{*}u=Ax+(-i)y+iz.$ (A.2)
Remark A.2.
The direct sum in (A.1) is not necessarily orthogonal.
Proposition A.3.
Let $A$ be a densely defined, closed, symmetric operator in some Hilbert space $H$ with deficiency indices equal $n_{\pm}(A)=1$. All self-adjoint extensions $A_{\theta}$ of $A$ may be parametrized by a real parameter $\theta\in[0,2\pi)$ where
$\begin{split}D(A_{\theta})&=\{x+c\phi_{+}+\zeta e^{i\theta}\phi_{-}:x\in D(A),\zeta\in\mathbb{C}\},\\ A_{\theta}(x+\zeta\phi_{+}+\zeta e^{i\theta}\phi_{-})&=Ax+i\zeta\phi_{+}-i\zeta e^{i\theta}\phi_{-},\end{split}$
with $A^{*}\phi_{\pm}=\pm i\phi_{\pm}$, and $\|\phi_{+}\|=\|\phi_{-}\|$.
By completeness in the exposition, we establish a extension theory for the Laplacian operator
$-\Delta=\Big{(}-\frac{d^{2}}{dx^{2}}\Big{)}_{\mathbb{e}\in\mathbb{E}}$ (A.3)
on a tadpole graph $\mathcal{G}$ with structure represented by the set $\mathbb{E}=(-L,L)\cup(L,+\infty)$ and a single vertex $\nu=L$, such that it induces the one-parameter family of self-adjoint operators $(-\Delta,D_{Z})_{Z\in\mathbb{R}}$ in (5.1) and (1.7). With these notations we have the following result.
Theorem A.4.
The Schrödinger type operator $-\Delta$ in (A.3) on $L^{2}(\mathcal{G})$, with domain
$D(-\Delta)$ =$=\{(\phi,\psi)\in H^{2}(\mathcal{G}):\phi(-L)=\phi(L)=\psi(L)=0,\;\;\phi^{\prime}(L)-\phi^{\prime}(-L)=\psi^{\prime}(L)\}$ (A.4)
is a densely defined symmetric operator with deficiency indices $n_{\pm}(-\Delta)=1$. Therefore, $(-\Delta,D(-\Delta))$ has a one-parameter family of self-adjoint extensions defined by $(-\Delta,D_{Z})_{Z\in\mathbb{R}}$ with
$D_{Z}=\{(f,g)\in H^{2}(\mathcal{G}):f(L)=f(-L)=g(L),\;\text{and},\;f^{\prime}(L)-f^{\prime}(-L)=g^{\prime}(L)+Zg(L)\}.$ (A.5)
Proof.
The symmetric property of $(-\Delta,D(-\Delta))$ is immediate. Since, $C_{c}^{\infty}(-L,L)\bigoplus C_{c}^{\infty}(L,+\infty)\subset D(-\Delta)$ we obtain the density property of $D(-\Delta)$ in $L^{2}(\mathcal{G})$. Now, it is not difficult to see that the adjoint operator $(-\Delta^{*},D(-\Delta^{*}))$ of $(-\Delta,D(-\Delta))$ is given by
$-\Delta^{*}=-\Delta,\quad D(-\Delta^{*})=\{(u,v)\in H^{2}(\mathcal{G}):u(-L)=u(L)=v(L)\}.$ (A.6)
Next, the deficiency subspaces $\mathcal{D}_{\pm}=\ker(-\Delta^{*}\mp i)$ have dimension one. Indeed, by using the definition of $D(-\Delta^{*})$ in (A.6) we can see that $\mathcal{D}_{\pm}=\mathrm{span}\{(f_{\pm i},g_{\pm i})\}$ with
[…]
and,
[…]
Thus, by Proposition A.3, the formulae in (A.7)-(A.8), and the von-Neumann decomposition in Theorem A.1, we can deduce that $(-\Delta,D(-\Delta))$ has a one-parameter family of self-adjoint extensions $(-\Delta,D_{Z})_{Z\in\mathbb{R}}$ with $D_{Z}$ defined in (A.5). This finishes the proof. ∎
A.2 Morse index estimate
In the following we obtain a general estimate of the Morse index for the family of self-adjoint operators $(\mathcal{L}_{Z},D_{Z})$ in (5.1)-(1.7).
We start with the following proposition which provides a strategy for estimating the Morse index of self-adjoint extensions of a minimal symmetric operator (see [37], [44], Chapter X). We recall the notation $n_{\pm}(A)=\dim\ker(A^{*}\mp iI)$.
Proposition A.5.
Let $A$ be a densely defined lower semi-bounded symmetric operator (that is, $A\geq mI$) with finite deficiency indices, $n_{\pm}(A)=k<\infty$, in the Hilbert space ${H}$, and let $\widehat{A}$ be a self-adjoint extension of $A$. Then the spectrum of $\widehat{A}$ in $(-\infty,m)$ is discrete and consists of, at most, $k$ eigenvalues counting multiplicities.
Next, let $\mathcal{G}_{0}$ be the graph determined by the set $\mathbb{E}=(-L,0)\cup(0,L)\cup(L,+\infty)$ and $H^{2}(\mathcal{G}_{0})=H^{2}(-L,0)\oplus H^{2}(0,L)\oplus H^{2}(L,+\infty)$. For $\mathbb{U}=(u,w,v)\in H^{2}(\mathcal{G}_{0})$ we define the Laplacian operator $-\Delta_{0}$ on $\mathcal{G}_{0}$ in the natural form
$-\Delta_{0}\mathbb{U}=(-u^{\prime\prime},-w^{\prime\prime},-v^{\prime\prime}).$ (A.9)
With this notations we have the following results (via a translation, all the statements below can be considered for $\mathbb{E}_{1}=(-L,0)\cup(0,L)\cup(0,+\infty)$ with vertex of the graph at $\nu=0$).
Theorem A.6.
The Schrödinger type operator $-\Delta_{0}$ in (A.9) on $L^{2}(\mathcal{G}_{0})$, with domain
$D(-\Delta_{0})=\{(u,w,v)\in H^{2}(\mathcal{G}_{0}):$ $u(-L)=w(L)=v(L)=0,\;u(0-)=w(0+)=0,$ (A.10)
$u^{\prime}(0-)=w^{\prime}(0+),\;w^{\prime}(L-)-u^{\prime}(-L+)=v^{\prime}(L+)\},$
is a densely defined symmetric operator with deficiency indices $n_{\pm}(-\Delta_{0})=2$. Therefore, we obtain that $(-\Delta_{0},D(-\Delta_{0}))$ has a four-parameter family of self-adjoint extensions. In particular, we have the family a two-parameter $(-\Delta_{0},D_{\lambda,\gamma})_{(\lambda,\gamma)\in\mathbb{R}^{2}}$ with
$D_{\lambda,\gamma}=\{(u,w,v)\in H^{2}(\mathcal{G}_{0}):$ $\;u(0-)=w(0+),\;u(-L)=w(L)=v(L),\;w^{\prime}(0+)-u^{\prime}(0-)=\lambda w(0+),$ (A.11)
$w^{\prime}(L-)-u^{\prime}(-L+)=v^{\prime}(L+)+\gamma v(L)\}.$
Thus, for $\lambda=0$, the family $(-\Delta_{0},D_{0,\gamma})_{(0,\gamma)\in\mathbb{R}^{2}}$ can be identified with the family $(-\Delta,D_{\gamma})_{\gamma\in\mathbb{R}}$ defined in Theorem A.4. In other words, $(-\Delta,D_{\gamma})_{\gamma\in\mathbb{R}}$ belongs to the self-adjoint extensions of the symmetric operator $(-\Delta_{0},D(-\Delta_{0}))$.
Proof.
The symmetry property of $(-\Delta_{0},D(-\Delta_{0}))$ is immediate. The density property of $D(-\Delta_{0})$ in $L^{2}(\mathcal{G}_{0})$ follows from $C_{c}^{\infty}(-L,0)\bigoplus C_{c}^{\infty}(0,L)\bigoplus C_{c}^{\infty}(L,+\infty)\subset D(-\Delta_{0})$. Now, by following a similar analysis as in Proposition A.5 in [14], the adjoint operator $(-\Delta_{0}^{*},D(-\Delta_{0}^{*}))$ of $(-\Delta_{0},D(-\Delta_{0}))$ is given by
$-\Delta_{0}^{*}=-\Delta_{0},\quad D(-\Delta_{0}^{*})=\{(u,w,v)\in H^{2}(\mathcal{G}_{0}):u(0-)=w(0+),\;u(-L)=w(L)=v(L)\}.$ (A.12)
Next, the deficiency subspaces $\mathcal{D}_{\pm}=\ker(-\Delta_{0}^{*}\mp i)$ have dimension 2. Indeed, by using the definition of $D(-\Delta_{0}^{*})$ in (A.12) we can see that $\mathcal{D}_{\pm}=\mathrm{span}\{\Pi_{\pm i},\Theta_{\pm i}\}$ with
$\Pi_{-i}=\Big{(}\underset{\bf-L<x<0}{e^{\sqrt{i}(x+L)}},\underset{\bf 0<x<L}{e^{-\sqrt{i}(x-L)}},\underset{\bf L<x}{e^{\sqrt{i}(x-L)}}\Big{)},\;\;\Theta_{-i}=\Big{(}\underset{\bf-L<x<0}{e^{-\sqrt{i}(x+L)}},\underset{\bf 0<x<L}{e^{\sqrt{i}(x-L)}},\underset{\bf L<x}{e^{\sqrt{i}(x-L)}}\Big{)},\;\;\mathrm{Im}(\sqrt{i})<0,$
and
$\Pi_{+i}=\Big{(}\underset{\bf-L<x<0}{e^{\sqrt{-i}(x+L)}},\underset{\bf 0<x<L}{e^{-\sqrt{-i}(x-L)}},\underset{\bf L<x}{e^{\sqrt{-i}(x-L)}}\Big{)},\;\;\Theta_{+i}=\Big{(}\underset{\bf-L<x<0}{e^{-\sqrt{-i}(x+L)}},\underset{\bf 0<x<L}{e^{\sqrt{-i}(x-L)}},\underset{\bf L<x}{e^{\sqrt{-i}(x-L)}}\Big{)},\;\;\mathrm{Im}(\sqrt{-i})>0.$
Thus, from the Krein and von Neumann theory, we deduce that $(-\Delta_{0},D(-\Delta_{0}))$ has a four-parameter family of self-adjoint extensions, in particular, we obtain one family a two-parameter of self-adjoint extensions $(-\Delta_{0},D_{\lambda,\gamma})_{(\lambda,\gamma)\in\mathbb{R}^{2}}$ determined by $D_{\lambda,\gamma}$ in (A.11). The statement about $\lambda=0$ is immediate. This finishes the proof. ∎
The domains $D_{\lambda,\gamma}$ can be seen as $\delta$-interaction type domains with “deltas” located at the points $x=0,L$. The introduction of the graph $\mathcal{G}_{0}$ is a main piece for obtaining a Morse index estimate of $(\mathcal{L}_{Z},D_{Z})$ in (5.1) (see Theorem 5.6 and Theorem A.7 below). We note that this strategy of splitting the graph $\mathcal{G}$ at $x=0$ is due to the fact that the positive single-lobe profile $\phi_{1}$ on $[-L,L]$ satisfies $\phi_{1}^{\prime}(0)=0$ and so Proposition A.5 can be used.
Theorem A.7.
Let $(\phi_{1},\phi_{2})\in D_{Z}$ be a positive single-lobe kink state for system (1.6). Then, for $(\mathcal{L}_{Z},D_{Z})$ defined in (5.1) we have the Morse index estimate, $n(\mathcal{L}_{Z})\leqq 2$.
Proof.
The idea of the proof is to split the operator $\mathcal{L}_{Z}$ on a convenient graph and to apply the extension theory of symmetric operators. Without loss of generality, we consider $c_{1}=c_{2}=1$. Then, for the metric graph $\mathcal{G}_{0}$ defined above, we consider the diagonal-matrix $\mathcal{M}=\mathrm{diag}(M_{i})_{1\leqq i\leqq 3}$ with
$M_{1}u(x)$ $=(-\partial_{x}^{2}+\cos(\phi_{1}))\,u(x),$ $x\in(-L,0)$ (A.13)
$M_{2}w(x)$ $=(-\partial_{x}^{2}+\cos(\phi_{1}))\,w(x),$ $x\in(0,L)$
$M_{3}v(x)$ $=(-\partial_{x}^{2}+\cos(\phi_{2}))\,v(x),$ $x>L.$
Then, $(\mathcal{M},D(\mathcal{M}))$, with $D(\mathcal{M})=D(-\Delta_{0})$ in (A.10) is a symmetric operator and the deficiency indices satisfy $n_{\pm}(\mathcal{M})=n_{\pm}(-\Delta_{0})=2$ (because $\mathcal{M}-\Delta_{0}$ is a bounded operator on $L^{2}(\mathcal{G}_{0})$). Moreover, there is a two-parameter family of self-adjoint extensions $(\mathcal{M}_{\lambda,\gamma},D(\mathcal{M}_{\lambda,\gamma}))_{(\lambda,\gamma)\in\mathbb{R}^{2}}$ associated to $(\mathcal{M},D(\mathcal{M}))$ with $\mathcal{M}_{\lambda,\gamma}=\mathcal{M}$ and $D(\mathcal{M}_{\lambda,\gamma})=D_{\lambda,\gamma}$ defined in (A.11). Next, we see $\mathcal{M}\geqq 0$. Indeed, for ${\bf{U}}=(u,w,v)\in D(-\Delta_{0})$ and from the (well-defined) formulae
$M_{1}u$ $=-\frac{1}{\phi_{1}^{\prime}}\frac{d}{dx}\Big{[}(\phi_{1}^{\prime})^{2}\frac{d}{dx}\Big{(}\frac{u}{\phi_{1}^{\prime}}\Big{)}\Big{]},\quad$ $x\in(-L,0),$ (A.14)
$M_{2}w$ $=-\frac{1}{\phi_{1}^{\prime}}\frac{d}{dx}\Big{[}(\phi_{1}^{\prime})^{2}\frac{d}{dx}\Big{(}\frac{w}{\phi_{1}^{\prime}}\Big{)}\Big{]},\quad$ $x\in(0,L),$
$M_{3}v$ $=-\frac{1}{\phi_{2}^{\prime}}\frac{d}{dx}\Big{[}(\phi_{2}^{\prime})^{2}\frac{d}{dx}\Big{(}\frac{v}{\phi_{2}^{\prime}}\Big{)}\Big{]},\quad$ $x>L,$
we obtain via integration by parts the relation
[…]
where the $I_{i}$’s are the remainders defined by
$I_{1}$ $=-\frac{u}{\phi_{1}^{\prime}}(\phi_{1}^{\prime})^{2}\frac{d}{dx}\Big{(}\frac{u}{\phi_{1}^{\prime}}\Big{)}\Big{|}_{-L}^{0},$ (A.16)
$I_{2}$ $=-\frac{w}{\phi_{1}^{\prime}}(\phi_{1}^{\prime})^{2}\frac{d}{dx}\Big{(}\frac{w}{\phi_{1}^{\prime}}\Big{)}\Big{|}_{0}^{L},$
$I_{3}$ $=-\frac{v}{\phi_{2}^{\prime}}(\phi_{2}^{\prime})^{2}\frac{d}{dx}\Big{(}\frac{v}{\phi_{2}^{\prime}}\Big{)}\Big{|}_{L}^{\infty}.$
Next, we see $I_{i}=0$ for $i=1,2,3$. Indeed, since $u(0-)=u(-L)=0$, $\phi_{1}^{\prime}(-L)\neq 0$, $\phi_{1}^{\prime}(0)=0$ and $\phi_{1}^{\prime\prime}(0)\neq 0$, it follows from L’Hopital’s rule that
$I_{1}=\phi_{1}^{\prime\prime}(0)\lim_{\epsilon\to 0^{-}}\frac{u^{2}(\epsilon)}{\phi_{1}^{\prime}(\epsilon)}=0.$
Similarly, we get $I_{2}=0$. For $I_{3}$ we have from $v(L)=0$ and $\phi_{2}^{\prime}(L)\neq 0$,
$I_{3}=v(L)v^{\prime}(L)-v^{2}(L)\frac{\phi_{2}^{\prime\prime}(L)}{\phi_{2}^{\prime}(L)}=0.$
Due to Proposition A.5 and Theorem A.6, the family of self-adjoint extensions $(\mathcal{M}_{\lambda,\gamma},D(\mathcal{M}_{\lambda,\gamma}))_{(\lambda,\gamma)\in\mathbb{R}^{2}}$ satisfies $n(\mathcal{M}_{\lambda,\gamma})\leqq 2$ for every $(\lambda,\gamma)\in\mathbb{R}^{2}$, and for $\lambda=0$ we get $n(\mathcal{L}_{Z})\leqq 2$. This finishes the proof. ∎
A.3 Perron-Frobenius property
In what follows, we establish the Perron-Frobenius property for the family of self-adjoint operators (~$(\widetilde{\mathcal{L}},D_{\delta,\gamma})$, where
~$\widetilde{\mathcal{L}}=-c_{2}^{2}\partial_{x}^{2}+\cos(\psi_{\mathrm{even}}),$ (A.17)
$\psi_{\mathrm{even}}$ is the even-extension of the kink-profile $\psi_{a}(x)=4\arctan(e^{-\frac{1}{c_{2}}(x+a)})$, $x>0$, on all the line, and
$D_{\delta,\gamma}=\{f\in H^{2}(\mathbb{R}-\{0\})\cap H^{1}(\mathbb{R}):f^{\prime}(0+)-f^{\prime}(0-)=\gamma f(0)\}$ (A.18)
for any $\gamma\in\mathbb{R}$. So, we start with the following remark: by Weyl’s essential spectrum theorem (cf. Reed and Simon [45]), we have that the essential spectrum, $\sigma_{\mathrm{ess}}(\widetilde{\mathcal{L}})$, of $\widetilde{\mathcal{L}}$ satisfies $\sigma_{\mathrm{ess}}(\widetilde{\mathcal{L}})=[1,+\infty)$ for any $\gamma$.
Theorem A.8 (Perron-Frobenius property).
Consider the family of self-adjoint operators defined in (A.17)-(A.18), (~$(\widetilde{\mathcal{L}},D_{\delta,\gamma})_{\gamma\in\mathbb{R}}$. For $\gamma$ fixed, assume that $\beta=\inf\sigma(\widetilde{\mathcal{L}})<1$ is the smallest eigenvalue. Then, $\beta$ is simple, and its corresponding eigenfunction $\zeta_{\beta}$ is positive (after replacing $\zeta_{\beta}$ by $-\zeta_{\beta}$ if necessary) and even.
Proof.
This result follows by a slight twist of standard abstract Perron-Frobenius arguments (see Proposition 2 in Albert et al. [3]). Without loss of generality we consider $c_{2}=1$. The basic point in the analysis is to show that the Laplacian operator $-\Delta_{\gamma}\equiv-\frac{d^{2}}{dx^{2}}$ on the domain $D_{\delta,\gamma}$ has its resolvent $R_{\mu}=(-\Delta_{\gamma}+\mu)^{-1}$ represented by a positive kernel for some $\mu>0$ sufficiently large. Namely, for $f\in L^{2}(\mathbb{R})$
[…]
with $K(x,y)>0$ for all $x,y\in\mathbb{R}$. By convenience of the reader we show this main point, the remainder of the proof follows the same strategy as in [3]. Thus, for $\gamma$ fixed, let $\mu>0$ be sufficiently large (with $-2\sqrt{\mu}<\gamma$ in the case $\gamma<0$), then from the Krein formula (see Theorem 3.1.2 in [4]) we obtain
$K(x,y)=\frac{1}{2\sqrt{\mu}}\Big{[}e^{-\sqrt{\mu}|x-y|}-\frac{\gamma}{\gamma+2\sqrt{\mu}}e^{-\sqrt{\mu}(|x|+|y|)}\Big{]}.$
Moreover, for every $x$ fixed, $K(x,\cdot)\in L^{2}(\mathbb{R})$. Thus, the existence of the integral above is guaranteed by Holder’s inequality. Now, since $K(x,y)=K(y,x)$, it is sufficient to show that $K(x,y)>0$ in the following cases.
  1. (1)
    Let $x>0$ and $y>0$ or $x<0$ and $y<0$: for $\gamma\geqq 0$, we obtain from $\frac{\gamma}{\gamma+2\sqrt{\mu}}<1$ and $|x-y|\leqq|x|+|y|$, that $K(x,y)>0$. For $\gamma<0$ and $-2\sqrt{\mu}<\gamma$, follows immediately $K(x,y)>0$.
  2. (2)
    Let $x>0$ and $y<0$: in this case,
    $K(x,y)=\frac{1}{\gamma+2\sqrt{\mu}}e^{-\sqrt{\mu}(x-y)}>0$
    for any value of $\gamma$ (where again $-2\sqrt{\mu}<\gamma$ in the case $\gamma<0$) .